### v2

# Optimal Projection for 3D Gaussian Splatting

Abstract
3D Gaussian Splatting has garnered extensive attention and application in real-time neural rendering. Concurrently, concerns have been raised about the limitations of this technology in aspects such as point cloud storage, performance , and robustness in sparse viewpoints , leading to various improvements. However, there has been a notable lack of attention to the projection errors introduced by the local affine approximation inherent in the splatting itself, and the consequential impact of these errors on the quality of photo-realistic rendering. This paper addresses the projection error function of 3D Gaussian Splatting, commencing with the residual error from the first-order Taylor expansion of the projection function $\phi$. The analysis establishes a correlation between the error and the Gaussian mean position. Subsequently, leveraging function optimization theory, this paper analyzes the function’s minima to provide an optimal projection strategy for Gaussian Splatting referred to Optimal Gaussian Splatting. Experimental validation further confirms that this projection methodology reduces artifacts, resulting in a more convincingly realistic rendering.
Keywords:
3D deep learning, view synthesis, radiance fields, 3D gaussians, real-time rendering, error analysis
Figure 1: By minimizing the projection error through error analysis, we have achieved an improvement in the rendering image quality compared to the original 3D-GS [16].
1 Introduction
Novel view synthesis aims to generate new views of a 3D scene or object by interpolating from a set of images with known camera parameters, which is a longstanding challenge in computer vision and a prerequisite for many augmented reality (AR) and virtual reality (VR) applications. From traditional techniques such as Structure-From-Motion (SFM) [31] or Image-Based Rendering (IBR), to neural network-based scene representations [15, 30, 8, 24, 29, 32], various approaches have struggled to address the problem of synthesizing a new view from captured images. Significant progress in this field emerged with the advent of Neural Radiance Fields [25] based on MLPs (Multilayer Perceptrons). Subsequent works have extensively utilized neural radiance fields, including MipNeRF [1] to address aliasing and over blurriness in Neural Radiance Fields, and DietNeRF [14], RegNeRF [28], FreeNeRF [40], etc., to enhance the robustness of NeRF in sparse-view scenarios. NeRF++ [42] analyzes the success factors of NeRF and introduces the concept of Foreground-Background NeRF. Barron et al. introduced MipNeRF360 [2] as an extension of MipNeRF [1] to address the issue of generating low-quality renderings for unbounded scenes in NeRF. However, these methods rely on MLPs with a relatively large number of layers, leading to lengthy training and rendering times, which still fall significantly short of real-time rendering. As a result, many methods resort to auxiliary data structures for acceleration [26, 37, 7, 36, 39]. However, due to the implicit representation of scenes and the dense sampling of points along rays, they still face challenges in achieving real-time performance.
Recently, 3D Gaussian Splatting (3D-GS) [16] has departed from the implicit scene representation using MLPs and instead opted for an explicit representation using Gaussian functions. This approach avoids the need for sampling points during ray marching for rendering integration, replacing it with finite Gaussian splatting into the image space and achieving real-time rendering. Subsequently, a cascade of works and applications has emerged building upon the foundation of 3D-GS. SuGaR [11] is used for efficient 3D mesh reconstruction, aligning surfaces by regularizing geometry through the alignment of Gaussians. Recently, there have been efforts to enhance the robustness in sparse-view scenarios [38], performance, and storage efficiency [27] of 3D Gaussian Splatting. However, these improvement efforts have not specifically addressed the errors associated with Gaussian projection itself. A potential method that involves improvements to Gaussian functions is [41]. However, it introduces a Gaussian filter without investigating the projection function and projection errors. The approach still relies on projecting onto the $z=1$ plane.
Similar to the NeRF [25], the input to 3D Gaussian Splatting (3D-GS) consists of a set of images, together with the corresponding cameras calibrated by SfM [31]. However, in contrast to NeRF, 3D Gaussian Splatting takes the sparse point cloud generated during the SfM calibration process as input. From these points, it constructs a set of 3D Gaussians as graphical primitives explicitly representing the scene. Each 3D Gaussian is characterized by its position (mean) $\boldsymbol{\mu}$, covariance matrix $\boldsymbol{\Sigma}$, and opacity $\alpha$, and carries anisotropic spherical harmonics (SH) representing directional appearance component (color) of the radiance field, following standard practice [26, 7]. Subsequently, these 3D Gaussians are projected onto the image plane ($z=1$ plane) via a projection function denoted as $\phi$ for differentiable rasterization. This process, devoid of NeRF’s ray marching and points sampling for rendering integral, leads to a significant performance improvement compared to the NeRF based methods, achieving real-time rendering.
Unfortunately, Gaussian functions, despite retaining their Gaussian properties through convolution or affine transformation, may not necessarily maintain such characteristics after undergoing a projection transformation. Therefore, 3D Gaussian Splatting adopts a local affine approximation [44], specifically approximating the projection function with the first two terms of its Taylor expansion. Nevertheless, approximations introduce corresponding errors, and the local affine approximation similarly contributes to these errors which may lead to artifacts in the rendered images. This paper exploits the relationship between the error of 3D Gaussian Splatting and the Gaussian mean through the analysis of the Taylor remainder term. Additionally, it identifies the circumstances under which the error is minimized by determining the extremum of the error function.
Finally, based on the extremum analysis of the error function, we propose an optimal projection method. Specifically, we project along the direction from the Gaussian mean to the camera center, where the projection plane is tangent to the line connecting the Gaussian mean and the camera center. Further validation has been conducted on diverse datasets, comparing the proposed projection with the straightforward approach of projecting 3D Gaussians onto the $z=1$ plane. The results illustrate a reduced expected error, leading to real-time, realistic rendering, as shown in Figure 5 and Table 1.
To summarize, we provide the following contributions:
  • We analyze error of 3D Gaussian Splatting that may lead to artifacts and consequently degrades rendering quality. We have identified the correlation between this error and the Gaussian position.
  • We derive the mathematical expectation of this error function and analyze when this function takes extrema through methods of function optimization.
  • Based on the conclusion regarding the extremum of the expectation of this error function, we propose a new optimal projection method. Instead of naively projecting all Gaussians onto the same plane, we now adopt different tangent plane projections based on the directions from each Gaussian mean to the camera center. Through experiments, we further validated the superiority of this method. In various datasets and scenes, it consistently achieves better overall performance compared to the original projection method.
2 Related Work
To begin with, we will provide a succinct overview of traditional approaches in novel view synthesis and scene reconstruction. Subsequently, we will delve into works related to implicit scene representation using Neural Radiance Fields and explicit scene representation through 3D Gaussian Splatting.
2.1 Traditional Scene Reconstruction and Rendering
Early novel-view synthesis methods were based on light fields  [10, 19], transitioning from densely sampled to unstructured capture  [3]. Structure-from-Motion (SfM) [33] introduced a new era using photo collections for view synthesis. Multi-view stereo (MVS) and subsequent algorithms  [9, 5] blended input images but faced challenges. Recent neural rendering algorithms [18] have overcome these issues, outperforming traditional methods.
2.2 Neural Radiance Fields
The Neural Radiance Field (NeRF) [25] stands out as a successful neural rendering method based on MLPs, primarily owing to its encoding of position and direction. This encoding allows for effective reconstruction of high-frequency information in scenes. Notable improvements on this encoding have been made by MipNeRF [1], NeRF-W [22], FreeNeRF [40], and Instant NGP [26]. These enhancements enable the handling of multi-resolution image inputs, multi-illumination with occlusion image inputs, sparse-view inputs, and achieve nearly real-time rendering capabilities, respectively. Importantly, these methods thoroughly exploit the intrinsic capabilities of NeRF as an implicit scene representation without introducing additional model priors. However, these methods are constrained by the dense sampling of rays, preventing them from achieving real-time performance. In contrast, the recent 3D Gaussian Splatting [16] achieves real-time rendering by breaking away from MLPs and ray sampling constraints.
2.3 3D Gaussian Splatting
3D Gaussian Splatting [16] utilizes Structure-from-Motion (SFM) [31] point clouds as input to construct an explicit scene representation using Gaussian functions. Subsequently, these Gaussian functions undergo a series of coordinate transformations and are projected onto the image plane at $z=1$, where rasterization takes place.
The significant breakthrough in performance is achieved by eschewing MLPs and ray sampling in favor of Gaussian functions and Gaussian Splatting. This shift has garnered widespread attention in the industry, leading to various works building upon this method. Gaussian Splatting is applied in SLAM (Simultaneous Localization and Mapping) systems [23]. Additionally, there is scene editing [6] and segmentation [4] based on 3D Gaussian.
There are also efforts to integrate 3D Gaussian Splatting with the popular diffusion models [13, 34, 35] in the AIGC domain [20]. The diffusion model is employed to provide geometric priors for 3D Gaussian Splatting, correcting inconsistencies in multi-view geometry. Another method involves tracking dynamic Gaussians through continuous dynamic view synthesis [21]. However, these enhancement efforts have relatively less emphasis on the 3D Gaussian Splatting method itself and often involve the use of additional models or priors to assist 3D-GS.
This paper aims to explore the potential of 3D Gaussian Splatting by analyzing the errors that may arise during the projection process. The analysis will delve into the factors contributing to these errors and propose methods to mitigate and reduce them.
3 Preliminaries
(a) 0.095
(b) 0.95
Figure 2: Visualizations of the 3D Gaussian Splatting error function $\epsilon\left(\theta_{\mu},\phi_{\mu}\right)$. The two subplots from left to right respectively depict the function graphs within the domains of $\left\{\left(\theta_{\mu},\phi_{\mu}\right)\mid\theta\in\left[-\pi/4\cdot 0.095,\pi/4\cdot 0.095\right]~\land~\phi\in\left[-\pi/4\cdot 0.095,\pi/4\cdot 0.095\right]\right\}$ and $\left\{\left(\theta_{\mu},\phi_{\mu}\right)\mid\theta\in\left[-\pi/4\cdot 0.95,\pi/4\cdot 0.95\right]~\land~\phi\in\left[-\pi/4\cdot 0.95,\pi/4\cdot 0.95\right]\right\}$ where $\pi/4$ corresponds to the integration region of the mathematical expectation. These representations are employed for a better understanding of the nature of this error function.
In the world coordinate system, let there be a 3D Gaussian function $G$, characterized by its mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$, expressed as:
$G_{\mathbf{\mu},\boldsymbol{\Sigma}}\left(\mathbf{x}\right)=\exp\left(-1/2\left(\mathbf{x}-\boldsymbol{\mu}\right)^{\top}\boldsymbol{\Sigma}^{-1}\left(\mathbf{x}-\boldsymbol{\mu}\right)\right)\text{.}$ (1)
This Gaussian is multiplied by $\alpha$ in blending process.
To project the Gaussian onto the image plane, the initial step in 3D Gaussian Splatting involves an affine transformation of this Gaussian function from the world coordinate system to the camera coordinate system via a viewing transformation matrix $\mathbf{W}$. The transformed Gaussian function is as follows:
$\begin{split}G\left(\mathbf{x}\right)&=\exp\left(-1/2\left(\mathbf{x}-\boldsymbol{\mu}\right)^{\top}\boldsymbol{\Sigma}^{-1}\left(\mathbf{x}-\boldsymbol{\mu}\right)\right)\\ &=\exp\left(-1/2\left(\mathbf{W}\mathbf{x}-\mathbf{W}\boldsymbol{\mu}\right)^{\top}\left(\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}\right)^{-1}\left(\mathbf{W}\mathbf{x}-\mathbf{W}\boldsymbol{\mu}\right)\right)\end{split}$ (2)
which can also be represented by the following function:
$\begin{split}G^{{}^{\prime}}\left(\mathbf{x^{{}^{\prime}}}\right)&=\exp\left(-1/2\left(\mathbf{x^{{}^{\prime}}}-\boldsymbol{\mu^{{}^{\prime}}}\right)^{\top}{\boldsymbol{\Sigma}^{{}^{\prime}}}^{-1}\left(\mathbf{x^{{}^{\prime}}}-\boldsymbol{\mu^{{}^{\prime}}}\right)\right)\end{split}$ (3)
where $G^{{}^{\prime}}$, $\mathbf{x}^{{}^{\prime}}=\mathbf{W}\mathbf{x}$, $\boldsymbol{\mu}^{{}^{\prime}}=\mathbf{W}\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}^{{}^{\prime}}=\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}$ represent the 3D Gaussian function, point, mean and covariance matrix in the camera coordinate system, respectively.
Subsequently, it is necessary to project the Gaussian function in the camera coordinate system onto the $z=1$ plane. The projection plane is formulated as:
$\mathbf{x_{0}}^{\top}\cdot\left(\mathbf{x}^{{}^{\prime}}-\mathbf{x_{0}}\right)=0$ (4)
where $\mathbf{x_{0}}=\left[\begin{matrix}0,0,1\end{matrix}\right]^{\top}$ represents the projection of the camera coordinate system origin onto this plane. According to the projection plane equation (Equation 4), the projection function $\phi$ is obtained as:
$\phi\left(\mathbf{x}^{{}^{\prime}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{0}}^{T}\mathbf{x}^{{}^{\prime}}\right)^{-1}\left(\mathbf{x_{0}}^{T}\mathbf{x_{0}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{0}}^{T}\mathbf{x}^{{}^{\prime}}\right)^{-1}.$ (5)
Expanding this projection function to the first order using Taylor series yields:
$\begin{split}\phi\left(\mathbf{x}^{{}^{\prime}}\right)&=\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)+\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)+R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)\\ &\approx\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)+\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)\end{split}$ (6)
where $\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$ is the Jacobian of the affine approximation of the projective transformation, denoted as $\mathbf{J}$. Applying this local affine approximation by neglecting the Taylor remainder term allows to derive the 2D Gaussian function $G_{\text{2D}}$ projected onto the $z=1$ plane similar to Equation 2:
$\begin{split}G_{\text{2D}}\left(\mathbf{x}^{{}^{\prime}}\right)&=\exp\left(-1/2\left(\mathbf{J}\mathbf{x}^{{}^{\prime}}-\mathbf{J}\boldsymbol{\mu}^{{}^{\prime}}\right)^{\top}\left(\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}\right)^{-1}\left(\mathbf{J}\mathbf{x}^{{}^{\prime}}-\mathbf{J}\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\\ &\approx\exp\left(-1/2\left(\phi\left(\mathbf{x}^{{}^{\prime}}\right)-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)^{\top}\left(\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}\right)^{-1}\left(\phi\left(\mathbf{x}^{{}^{\prime}}\right)-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\right).\end{split}$ (7)
Note that since the rank of the matrix $\mathbf{J}$ is 2, the inverse of $\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}$ is, in fact, the inverse of the covariance matrix of the 2D Gaussian which means skipping the third row and column of this matrix. Similarly, this function can be expressed in another form:
$G_{\text{2D}}\left(\mathbf{x_{\text{2D}}}\right)=\exp\left(-1/2\left(\mathbf{x_{\text{2D}}}-\boldsymbol{\mu_{\text{2D}}}\right)^{\top}\boldsymbol{\Sigma}_{\text{2D}}^{-1}\left(\mathbf{x_{\text{2D}}}-\boldsymbol{\mu_{\text{2D}}}\right)\right)$ (8)
where $G_{\text{2D}}$, $\mathbf{x}_{\text{2D}}=\phi\left(\mathbf{x}^{{}^{\prime}}\right)$, $\boldsymbol{\mu}_{\text{2D}}=\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$ and $\boldsymbol{\Sigma}_{\text{2D}}=\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}$ represent the 2D Gaussian function, point, mean and $\times 2$ 2 covariance matrix in the image coordinate system, respectively.
After the aforementioned steps, 3D Gaussian Splatting projects the 3D Gaussian from the world coordinate system onto the image plane. Subsequently, rasterization is performed on the image plane to obtain the rendered image.
4 Local Affine Approximation Error
From Equation 6 and Equation 7 we observe that 3D Gaussian Splatting introduces an approximation during the projection transformation, i.e. the local affine approximation. The rasterization uses a 2D Gaussian function that is not the true projection function of the 3D Gaussian. The error introduced by this approximation can be characterized by the Taylor remainder term in Equation 6:
$\begin{split}&R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)=\phi\left(\mathbf{x}^{{}^{\prime}}\right)-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)-\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)\end{split}$ (9)
where
$\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\mathbb{I}\otimes\left(\mathbf{x_{0}}^{T}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}-\mathbf{x_{0}}\left(\mathbf{x_{0}}^{T}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{T}\mathbf{x_{0}}\right)^{-1}{\boldsymbol{\mu}^{{}^{\prime}}}^{T}$ (10)
and $\mathbb{I}$ represents the identity matrix; $\otimes$ denotes the multiplication of a matrix and a scalar. It is not difficult to observe that this Taylor remainder term is a three-dimensional vector related to the random variable $\mathbf{x}^{{}^{\prime}}$ and the mean $\boldsymbol{\mu}^{{}^{\prime}}$ of the Gaussian function. Therefore, by computing the square of the Frobenius norm of this vector $\|R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)\|_{F}^{2}$ and taking the mathematical expectation of this norm function with respect to the random variable $\mathbf{x}^{{}^{\prime}}$, we ultimately obtain an error function that depends solely on the Gaussian mean $\boldsymbol{\mu}^{{}^{\prime}}$:
$\epsilon\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\int\limits_{\mathbf{x}^{{}^{\prime}}\in\mathcal{X}^{{}^{\prime}}}\|R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)\|_{F}^{2}\mathrm{d}\mathbf{x}^{{}^{\prime}}\text{.}$ (11)
Before seeking the extremum of the error function, let’s simplify the integral expression. First, we simplify $\mathbf{x}^{{}^{\prime}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$ to unit vectors projected onto the unit sphere centered at the camera center (hereafter referred to as the unit sphere) by proving that the composition of transformations $\left(\phi\circ\pi\right)$, projecting a point in three-dimensional space onto the unit sphere $\pi$ and subsequently onto the projection plane $\phi$, is equivalent to directly projecting a point in three-dimensional space onto the projection plane. This simplification ensures that the entire error function involves only operations between three unit space vectors. The proof proceeds as follows:
$\begin{split}\left(\phi\circ\pi\right)\left({\mathbf{x}^{{}^{\prime}}}\right)&=\phi\left({\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\right)\\ &={\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\left(\mathbf{x_{0}}^{\top}\left({\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\right)\right)^{-1}\\ &={\mathbf{x}^{{}^{\prime}}}\left(\mathbf{x_{0}}^{T}{\mathbf{x}^{{}^{\prime}}}\right)^{-1}\\ &=\phi\left({\mathbf{x}^{{}^{\prime}}}\right)\text{.}\end{split}$ (12)
The simplified error function $\epsilon$ involves three unit vectors, or equivalently, three points on the unit sphere: $\mathbf{x}^{{}^{\prime}}$, $\mathbf{x_{0}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$, as illustrated in Figure 3. Their spherical coordinates are given as follows:
$\mathbf{x_{0}}=\left[\begin{matrix}\sin{\left(\phi_{0}\right)}\cos{\left(\theta_{0}\right)}\\ -\sin{\left(\theta_{0}\right)}\\ \cos{\left(\phi_{0}\right)}\cos{\left(\theta_{0}\right)}\end{matrix}\right],\mathbf{x}^{{}^{\prime}}=\left[\begin{matrix}\sin{\left(\phi\right)}\cos{\left(\theta\right)}\\ -\sin{\left(\theta\right)}\\ \cos{\left(\phi\right)}\cos{\left(\theta\right)}\end{matrix}\right],\boldsymbol{\mu}^{{}^{\prime}}=\left[\begin{matrix}\sin{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}\\ -\sin{\left(\theta_{\mu}\right)}\\ \cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}\end{matrix}\right]$ (13)
where $\phi_{0}=0$ and $\theta_{0}=0$. Based on Equation 9, Equation 10, and Equation 13, the following derivations can be obtained:
$\mathbf{J}=\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\left[\begin{matrix}\frac{1}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}&0&-\frac{\sin{\left(\phi_{\mu}\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}\\ 0&\frac{1}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}&\frac{\sin{\left(\theta_{\mu}\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\theta_{\mu}\right)}}\\ 0&0&0\end{matrix}\right]$ (14)
$R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)=\left[\begin{matrix}-\frac{\sin{\left(\phi-\phi_{\mu}\right)}\cos{\left(\theta\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}+\tan{\left(\phi\right)}-\tan{\left(\phi_{\mu}\right)}\\ \frac{\sin{\left(\theta\right)}}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}-\frac{\sin{\left(\theta_{\mu}\right)}\cos{\left(\phi\right)}\cos{\left(\theta\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\theta_{\mu}\right)}}+\frac{\tan{\left(\theta_{\mu}\right)}}{\cos{\left(\phi_{\mu}\right)}}-\frac{\tan{\left(\theta\right)}}{\cos{\left(\phi\right)}}\\ 0\end{matrix}\right]\text{.}$ (15)
The size of the integration region is related to the covariance of the Gaussian function. Since the current discussion specifically focuses on the impact of the mean on the error rather than the variance, let’s assume the integration region to be $\left\{\mathbf{x}^{{}^{\prime}}\mid\theta\in\left[-\pi/4+\theta_{\mu},\pi/4+\theta_{\mu}\right]~\land~\phi\in\left[-\pi/4+\phi_{\mu},\pi/4+\phi_{\mu}\right]\right\}$. Then, substituting Equation 15 into Equation 11, we can derive:
$\begin{split}\epsilon\left(\theta_{\mu},\phi_{\mu}\right)&=\int_{-\pi/4+\theta_{\mu}}^{\pi/4+\theta_{\mu}}\int_{-\pi/4+\phi_{\mu}}^{\pi/4+\phi_{\mu}}\left(-\frac{\sin{\left(\phi-\phi_{\mu}\right)}\cos{\left(\theta\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}+\tan{\left(\phi\right)}-\tan{\left(\phi_{\mu}\right)}\right)^{2}+\\ &\left(\frac{\sin{\left(\theta\right)}}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}-\frac{\sin{\left(\theta_{\mu}\right)}\cos{\left(\phi\right)}\cos{\left(\theta\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\theta_{\mu}\right)}}+\frac{\tan{\left(\theta_{\mu}\right)}}{\cos{\left(\phi_{\mu}\right)}}-\frac{\tan{\left(\theta\right)}}{\cos{\left(\phi\right)}}\right)^{2}\mathrm{d}\theta\mathrm{d}\phi\text{.}\end{split}$ (16)
The function is partially differentiated with respect to $\theta_{\mu}$ and $\phi_{\mu}$. It is observed that at $\theta_{\mu}=\theta_{0}=0$ and $\phi_{\mu}=\phi_{0}=0$, the following holds:
$\frac{\partial\epsilon}{\partial\theta_{\mu}}\left(0,0\right)=0,\frac{\partial\epsilon}{\partial\phi_{\mu}}\left(0,0\right)=0\text{.}$ (17)
Therefore, this point is an extremum point of the function. Upon verification, it is confirmed that this point is the minimum value point of the function. We have visualized the error function, as shown in Figure 2. Through Subfigure 2a in Figure 2, it can be observed that this function is a concave function, and it attains its minimum value at (0,0)(0,0)( 0 , 0 ). Examining Subfigure 2b in Figure 2, it is evident that in the majority of the region close to the origin, the function values are not significantly different. However, as the function approaches the integration limits, the function values increase rapidly, resulting in a substantial difference between the maximum and minimum values.
In other words, in most cases, the error is small and does not significantly affect the quality of rendered images. This is why 3D Gaussian Splatting utilizes local affine approximation but still successfully reconstructs the scene to obtain a high-quality novel view image. Nevertheless, the naive projection of all gaussians onto the same plane $z=1$ in 3D Gaussian Splatting may lead to larger projection errors for gaussians farther from the plane center (the projection point of the camera center to the plane), causing artifacts, as shown in Figure 5.
Figure 3: Toy 2D visualizations of the projection transformation. $\mathbf{x}^{{}^{\prime}}$, $\mathbf{x_{0}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$ are all simplified unit vectors, representing the random variable, the tangency point of the projection plane and the unit sphere, and the Gaussian mean, respectively, denoted by green, purple, and blue arrows in the diagram. Due to the $3\sigma$ rule in the Gaussian distribution and the fact that actual rasterization does not sample all points, we represent the integration domain with a circular arc symmetric about the mean vector $\boldsymbol{\mu}^{{}^{\prime}}$.
5 Optimal Gaussian Splatting
Figure 4: Illustration of the rendering pipeline for Optimal Gaussian Splatting. The first step, similar to the original 3D-GS [16], involves transforming world coordinates into the camera coordinate system. Instead of projecting Gaussians onto the z=1 plane, each Gaussian is radially projected based on its mean along the line connecting it to the camera center, projecting onto a plane tangent to the unit sphere and perpendicular to this line. After projection, colors of points on the unit sphere are computed by alpha blending the 2D Gaussians on the tangent plane. Finally, for each pixel in the image, rays are cast onto the unit sphere to retrieve the color of that pixel, resulting in the rendered image.
Based on the analysis of the error function, in this section, we will introduce Optimal Gaussian Splatting shown in Figure 4 that exhibits smaller projection errors, resulting in higher-quality rendering, as shown in Figure 5 and Table 1.
5.1 Optimal Projection
Through the analysis of the error function in Section 4, we discovered that the error function attains its minimum value when the projection of the Gaussian mean on the plane coincides with the projection from the camera center to the plane. Therefore, in Optimal Gaussian Splatting, we employ an optimal projection method. Specifically, instead of naively projecting different Gaussians onto the same plane $z=1$, we adopt distinct projection planes for each Gaussian. These projection planes are determined by the tangent planes formed by the Gaussian mean and the line connecting it to the camera center. The specific projection plane is ormulated as:
$\mathbf{x^{\mu}}^{\top}\cdot\left(\mathbf{x}^{{}^{\prime}}-\mathbf{x^{\mu}}\right)=0$ (18)
where
$\mathbf{x^{\mu}}=\pi\left({\boldsymbol{\mu}^{{}^{\prime}}}\right)={\boldsymbol{\mu}^{{}^{\prime}}}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{\top}{\boldsymbol{\mu}^{{}^{\prime}}}\right)^{-1/2}$ (19)
represents the projection of the camera coordinate system origin onto this plane. According to the projection plane equation (Equation 18 and Equation19), the optimal projection function $\phi_{o}$ is obtained as:
$\phi^{\mu}\left(\mathbf{x}^{{}^{\prime}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x^{\mu}}^{T}\mathbf{x}^{{}^{\prime}}\right)^{-1}\left(\mathbf{x^{\mu}}^{T}\mathbf{x^{\mu}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x^{\mu}}^{T}\mathbf{x}^{{}^{\prime}}\right)^{-1}\text{.}$ (20)
The corresponding local affine approximation Jacobian matrix $\mathbf{J_{o}}$ is:
$\mathbf{J^{\mu}}=\frac{\partial\phi^{\mu}}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\mathbb{I}\otimes\left(\mathbf{x^{\mu}}^{T}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}-\mathbf{x^{\mu}}\left(\mathbf{x^{\mu}}^{T}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{T}\mathbf{x^{\mu}}\right)^{-1}{\boldsymbol{\mu}^{{}^{\prime}}}^{T}\text{.}$ (21)
When $\boldsymbol{\mu}^{{}^{\prime}}=\left[\begin{matrix}\mu_{x},\mu_{y},\mu_{z}\end{matrix}\right]^{\top}$, the specific form of this Jacobian matrix is:
$\mathbf{J^{\mu}}=\left[\begin{matrix}\frac{\mu_{y}^{2}+\mu_{z}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{x}\mu_{y}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{x}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\\ -\frac{\mu_{x}\mu_{y}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&\frac{\mu_{x}^{2}+\mu_{z}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{y}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\\ -\frac{\mu_{x}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{y}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&\frac{\mu_{x}^{2}+\mu_{y}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\end{matrix}\right]$ (22)
5.2 Unit Sphere Based Rasterizer
For Optimal Projection, we have employed a Unit Sphere Based Rasterizer for rasterization to generate images.
Through Optimal Projection, we obtain the projection of the three-dimensional Gaussians on the tangent plane of the unit sphere instead of obtaining the Gaussians in the image space. Therefore, we need to rasterize based on this unit sphere to generate images.
In specific terms, for a pixel $\left(u,v\right)$ on the image, similar to NeRF, we cast a ray. However, unlike NeRF, we do not involve the extensive sampling of points along the ray, which would significantly degrade performance. Instead, our focus is solely on determining which tangent-plane Gaussians the ray intersects on the unit sphere:
$\mathbf{x^{\mu}_{\text{2D}}}=\phi^{\mu}\left(\left[\begin{matrix}\left(u-c_{x}\right)/f_{x}\\ \left(v-c_{y}\right)/f_{y}\\ 1\end{matrix}\right]\right)$ (23)
We then query the function values of these Gaussians for alpha blending to obtain the color
$G^{\mu}_{\text{2D}}\left(\mathbf{x^{\mu}_{\text{2D}}}\right)=\exp\left(-1/2\left(\mathbf{x^{\mu}_{\text{2D}}}-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)^{\top}\left(\mathbf{J^{\mu}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J^{\mu}}^{\top}\right)^{-1}\left(\mathbf{x^{\mu}_{\text{2D}}}-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\right)$ (24)
$\alpha\left(u,v,\boldsymbol{\mu}\right)=\alpha_{\boldsymbol{\mu}}\cdot G^{\mu}_{\text{2D}}\left(\mathbf{x^{\mu}_{\text{2D}}}\right)$ (25)
$\mathbf{C}\left(u,v\right)=\sum_{z=z_{\text{near}}}^{z_{\text{far}}}{\text{SH}\left(\boldsymbol{\mu}_{z}\right)\alpha\left(u,v,\boldsymbol{\mu}_{z}\right)\prod_{k=z_{\text{near}}}^{z}\left(1-\alpha\left(u,v,\boldsymbol{\mu}_{k}\right)\right)}$ (26)
where the summation and product from $z_{\text{near}}$ to $z_{\text{far}}$ represent discretized depths sorted in ascending order from near to far for the depths of 3D Gaussians encountered by the ray, $\alpha_{\boldsymbol{\mu}}$ represents the transparency of the 3D Gaussian with its mean $\boldsymbol{\mu}$, $\boldsymbol{\mu}_{z}$ represents the mean of the Gaussian at depth $z$, and SH represents the spherical harmonics of this Gaussian.
It is easy to notice that both $\mathbf{J}^{\mu}$ in Equation 21 and $\mathbf{J}$ in Equation 10 are matrices with a rank of 2. Therefore, in fact, $\mathbf{J^{\mu}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J^{\mu}}^{\top}$ does not have an inverse. We perform a reversible transformation on $\mathbf{J}^{\mu}$ to make the third row zero. Consequently, $\mathbf{J^{\mu}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J^{\mu}}^{\top}$ becomes a matrix with both the third row and the third column being zero. This allows us to treat it similar to $\mathbf{J}$, skipping the third row and third column, and compute its inverse as if it were a full-rank $\times 2$ 2 matrix. Certainly, to ensure the equation still holds, this invertible matrix also needs to be used for left multiplication with $\mathbf{x^{\mu}_{\text{2D}}}$ and $\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$.
6 Experiments
Table 1: Quantitative evaluation of our method compared to previous work, computed over three datasets. The Mip-NeRF360 [2] dataset encompasses both Indoor and Outdoor scenarios, while Tanks. and Deep. indicates the average results over the Tank&\&&Templates [17] and Deep Blending [12] datasets, respectively. The term Average denotes the mean value of measurements across all scenarios in these datasets.
Dataset Average Mip-NeRF360 Tanks. and Deep.
[…]
Plenoxels [7] 22.77 0.666 0.457 23.08 0.625 0.463 22.07 0.757 0.444
INGP-Base [26] 24.49 0.699 0.373 25.30 0.671 0.371 22.67 0.760 0.377
INGP-Big [26] 24.93 0.724 0.336 25.59 0.699 0.331 23.44 0.781 0.347
M-NeRF360 [2] 27.11 0.803 0.241 27.69 0.792 0.237 25.81 0.830 0.251
3D-GS [16] 26.92 0.832 0.214 27.21 0.815 0.214 26.27 0.872 0.213
Ours 27.17 0.836 0.210 27.48 0.821 0.209 26.44 0.872 0.214
Figure 5: We show comparisons of ours to previous methods and the corresponding ground truth images from held-out test views. Due to space constraints, we have selected methods such as 3D-GS [16] and Mip-NeRF360 [2], which are comparable in terms of metrics, for qualitative comparisons. The scenes are, from the top down: Truck and Train from Tanks&Temples [17]; Playroom, DrJohnson from the Deep Blending dataset [12] and Room, Bonsai, Counter and Stump from Mip-NeRF360 dataset [2]; Non-obvious Differences in quality highlighted by arrows/insets.
To further validate the effectiveness of the Optimal Gaussian Splatting derived from theoretical analysis, we conducted a series of experiments, comparing it with the original 3D-GS [16] and some current state-of-the-art methods.
6.1 Implementation
We implemented Optimal Gaussian Splatting based on the PyTorch framework in 3D-GS [16] and wrote custom CUDA kernels for rasterization. We used the default parameters of 3D-GS to maintain consistency with the original 3D-GS and prevent other factors from introducing interference into the results.
6.2 Datasets
We tested our algorithm on a total of 13 real scenes which are the same as those used in the original 3D-GS [16] paper. In particular, we evaluated our approach on the complete set of scenes featured in Mip-NeRf360 [2], which currently represents the state-of-the-art in NeRF rendering quality. Additionally, we tested our method on two scenes from the Tanks & Temples dataset [17] and two scenes provided by Hedman et al.  [12]. The selected scenes exhibit diverse capture styles, encompassing both confined indoor environments and expansive, unbounded outdoor settings. In other words, our evaluation includes a wide range of scenes to assess the performance of our method comprehensively.
6.3 Comparisons
We compare our algorithm with 3D-GS [16] and three state-of-the-art NeRF-based approaches: Mip-NeRF360 [2], InstantNGP [26] and Plenoxels [7]. All methods were configured using the same settings as outlined in the 3D-GS paper to control variables, preventing the introduction of variables other than our Optimal Splatting method.
6.4 Results
Quantitative comparisons
We adopt a train/test split for datasets following the methodology proposed by Mip-NeRF360 [2]. Specifically, every 8th photo is reserved for testing, ensuring a consistent and meaningful basis for comparisons to generate error metrics. Standard metrics such as PSNR, LPIPS [43], and SSIM, widely used in the literature, are employed for evaluation; refer to Table 1 for details. It can be observed that, while the 3D-GS method significantly outperforms MipNeRF360 in rendering speed, its rendering quality slightly falls behind MipNeRF360, with a lower average PSNR across the 13 scenes. This might be attributed to MipNeRF360 not employing significant approximations throughout its entire pipeline. However, by replacing the original projection method of 3D-GS with our Optimal Projection, which incurs smaller errors, we not only maintain rendering speed far surpassing MipNeRF360 but also achieve a quality improvement. Furthermore, in terms of average results across the 13 scenes, our approach outperforms others, including 3D-GS and several NeRF-based methods, in PSNR, SSIM, and LPIPS.
Qualitative comparisons
Due to space constraints, for our qualitative experiments, we selected only two methods, 3D-GS and MipNeRF360, from Table 1, which have comparable metrics, as our baselines.
Please see Figure 5. From the comparative images, it is evident that our algorithm, compared to 3D-GS, reduces artifacts and defects while preserving more details. In the Truck scene, we address the absence of the steering wheel in the reconstruction by 3D-GS (indicated by the red arrow), while avoiding the pillar fragmentation introduced by MipNeRF360 (indicated by the blue arrow). Additionally, in the Train scene, we maintain the geometry of the train’s iron railing while avoiding the black floaters present in 3D-GS. In the PlayRoom scene, we achieve relatively better wire geometry. In the DrJohnson scene, our algorithm neither exhibits the fragmented appearance seen in 3D-GS nor the blurred details observed in MipNeRF360. In the Room and Bonsai scenes, we respectively have better stems in the flowerpot and improved details on the small white objects on the floor compared to 3D-GS. In the Counter scene, our floor texture is optimal. In the Stump scene, our tree branches do not fracture as observed in 3D-GS.
7 Discussion and Conclusions
Because Gaussian functions remain Gaussian after affine transformations, convolutions, and similar operations, but this is not necessarily the case for projection transformations, the 3D Gaussian Splatting process employs a Taylor first-order expansion to approximate this projection transformation as an affine transformation. However, such approximation inevitably introduces errors, which can result in artifacts, distortions, and inaccuracies in the learning comprehension of the scene as well as the generation of novel view images. This paper begins by examining the Taylor expansion remainder of the local affine approximation, obtaining the mathematical expectation of the error. This expectation is a function of the mean of the Gaussian function. By finding critical points, it is determined that the function attains its minimum when the projection of the mean of the Gaussian function on the plane coincides with the projection of the camera center on the plane. The visualization of this function is depicted in Figure 2. Based on the conclusions drawn from this theoretical analysis, we propose a more optimal projection method with reduced error, referred to as Optimal Gaussian Splatting. Subsequently, extensive experiments are conducted to further validate the superiority of this projection approach.
The original 3D-GS [16], despite exhibiting significantly superior performance compared to the NeRF-based MipNeRF360 [2] method, may suffer a slight compromise in quality due to errors introduced by projection, as opposed to MipNeRF360 [2]. In contrast, our Optimal Gaussian Splatting, by minimizing projection errors, unlocks a portion of the untapped potential of 3D-GS. This demonstrates that the representation of scenes using 3D-GS can achieve state-of-the-art (SOTA) levels in both quality and performance. Additionally, it is noteworthy that our entire error analysis process only considers the influence of Gaussian mean on the error function while treating the covariance as a constant. Hence, future research could explore optimizing 3D-GS further by considering the Gaussian covariance in the analysis.
Furthermore, our method, throughout its entire process, is only dependent on the transformation from image pixels to the cutting plane (corresponding to Equation 23) and its relationship with the image representation. Unlike the original 3D-GS, where the $z=1$ plane tends to favor perspective images due to the perspective projection viewport transformation involving the NDC’s $z=1$ plane, our approach offers more versatility. This makes our method potentially applicable to various image representations. However, this versatility may come with a slight performance trade-off or require appropriate code implementation to avoid extending training time due to the additional operation of pixel-to-tangent-plane transformation, as shown in Equation 23.
In summary, our work represents the first theoretical analysis that focuses on the projection error of 3D Gaussian Splatting. We have introduced a novel Splatting method Optimal Gaussian Splatting aimed at reducing projection errors, and through both theoretical analysis and empirical validation, demonstrated the superiority of this approach, as shown in Figure 1 and Figure 5.

### v4

@@
-Optimal Projection for 3D Gaussian Splatting
+https://letianhuang.github.io/op43dgs/
 Abstract
@@
 Concurrently, concerns have been raised about the limitations of this technology in aspects such as point cloud storage, performance, and robustness in sparse viewpoints, leading to various improvements.
-However, there has been a notable lack of attention to the projection errors introduced by the local affine approximation inherent in the splatting itself, and the consequential impact of these errors on the quality of photo-realistic rendering.
-This paper addresses the projection error function of 3D Gaussian Splatting, commencing with the residual error from the first-order Taylor expansion of the projection function $\phi$.
+However, there has been a notable lack of attention to the fundamental problem of projection errors introduced by the local affine approximation inherent in the splatting itself, and the consequential impact of these errors on the quality of photo-realistic rendering.
+This paper addresses the projection error function of 3D Gaussian Splatting, commencing with the residual error from the first-order Taylor expansion of the projection function.
 The analysis establishes a correlation between the error and the Gaussian mean position.
-Subsequently, leveraging function optimization theory, this paper analyzes the function’s minima to provide an optimal projection strategy for Gaussian Splatting referred to Optimal Gaussian Splatting.
+Subsequently, leveraging function optimization theory, this paper analyzes the function’s minima to provide an optimal projection strategy for Gaussian Splatting referred to Optimal Gaussian Splatting, which can accommodate a variety of camera models.
 Experimental validation further confirms that this projection methodology reduces artifacts, resulting in a more convincingly realistic rendering.
 Keywords:
-3D deep learning, view synthesis, radiance fields, 3D gaussians, real-time rendering, error analysis
+3D deep learning view synthesis radiance fields 3D gaussians real-time rendering error analysis
 Figure 1: By minimizing the projection error through error analysis, we have achieved an improvement in the rendering image quality compared to the original 3D-GS [17].
+These images are cropped from the complete images of the Train scene, in order to be highlighted.
 1 Introduction
-Novel view synthesis aims to generate new views of a 3D scene or object by interpolating from a set of images with known camera parameters, which is a longstanding challenge in computer vision and a prerequisite for many augmented reality (AR) and virtual reality (VR) applications.
-From traditional techniques such as Structure-From-Motion (SFM) [31] or Image-Based Rendering (IBR), to neural network-based scene representations [15, 30, 8, 24, 29, 32], various approaches have struggled to address the problem of synthesizing a new view from captured images.
-Significant progress in this field emerged with the advent of Neural Radiance Fields [25] based on MLPs (Multilayer Perceptrons).
-Subsequent works have extensively utilized neural radiance fields, including MipNeRF [1] to address aliasing and over blurriness in Neural Radiance Fields, and DietNeRF [14], RegNeRF [28], FreeNeRF [40], etc., to enhance the robustness of NeRF in sparse-view scenarios.
-NeRF++ [42] analyzes the success factors of NeRF and introduces the concept of Foreground-Background NeRF.
-Barron et al.
-introduced MipNeRF360 [2] as an extension of MipNeRF [1] to address the issue of generating low-quality renderings for unbounded scenes in NeRF.
-However, these methods rely on MLPs with a relatively large number of layers, leading to lengthy training and rendering times, which still fall significantly short of real-time rendering.
-As a result, many methods resort to auxiliary data structures for acceleration [26, 37, 7, 36, 39].
-However, due to the implicit representation of scenes and the dense sampling of points along rays, they still face challenges in achieving real-time performance.
-Recently, 3D Gaussian Splatting (3D-GS) [16] has departed from the implicit scene representation using MLPs and instead opted for an explicit representation using Gaussian functions.
-This approach avoids the need for sampling points during ray marching for rendering integration, replacing it with finite Gaussian splatting into the image space and achieving real-time rendering.
-Subsequently, a cascade of works and applications has emerged building upon the foundation of 3D-GS.
-SuGaR [11] is used for efficient 3D mesh reconstruction, aligning surfaces by regularizing geometry through the alignment of Gaussians.
-Recently, there have been efforts to enhance the robustness in sparse-view scenarios [38], performance, and storage efficiency [27] of 3D Gaussian Splatting.
-However, these improvement efforts have not specifically addressed the errors associated with Gaussian projection itself.
-A potential method that involves improvements to Gaussian functions is [41].
-However, it introduces a Gaussian filter without investigating the projection function and projection errors.
-The approach still relies on projecting onto the $z=1$ plane.
-Similar to the NeRF [25], the input to 3D Gaussian Splatting (3D-GS) consists of a set of images, together with the corresponding cameras calibrated by SfM [31].
-However, in contrast to NeRF, 3D Gaussian Splatting takes the sparse point cloud generated during the SfM calibration process as input.
-From these points, it constructs a set of 3D Gaussians as graphical primitives explicitly representing the scene.
-Each 3D Gaussian is characterized by its position (mean) $\boldsymbol{\mu}$, covariance matrix $\boldsymbol{\Sigma}$, and opacity $\alpha$, and carries anisotropic spherical harmonics (SH) representing directional appearance component (color) of the radiance field, following standard practice [26, 7].
-Subsequently, these 3D Gaussians are projected onto the image plane ($z=1$ plane) via a projection function denoted as $\phi$ for differentiable rasterization.
-This process, devoid of NeRF’s ray marching and points sampling for rendering integral, leads to a significant performance improvement compared to the NeRF based methods, achieving real-time rendering.
-Unfortunately, Gaussian functions, despite retaining their Gaussian properties through convolution or affine transformation, may not necessarily maintain such characteristics after undergoing a projection transformation.
-Therefore, 3D Gaussian Splatting adopts a local affine approximation [44], specifically approximating the projection function with the first two terms of its Taylor expansion.
-Nevertheless, approximations introduce corresponding errors, and the local affine approximation similarly contributes to these errors which may lead to artifacts in the rendered images.
-This paper exploits the relationship between the error of 3D Gaussian Splatting and the Gaussian mean through the analysis of the Taylor remainder term.
+Reconstructing 3D scenes from 2D images and synthesizing novel views has been a critical task in computer vision and graphics.
+Recently, the trend of novel view synthesis is spearheaded by Neural Radiance Fields (NeRFs) [27] and its variants [1, 2], which achieves photo-realistic rendering quality.
+NeRF represents scenes with Multi-Layer Perceptron (MLP) and renders novel views using volumetric ray-marching.
+However, these MLP-based methods contain a relatively large number of layers, leading to lengthy training and rendering times, which still fall significantly short of real-time rendering.
+To accelerate rendering, many methods resort to auxiliary data structures, such as hash grids [28], points or voxel [39, 9, 38, 41].
+However, due to the dense sampling of points along rays, they still face challenges in achieving real-time performance.
+With a substantial demand for high-speed and photo-realistic rendering effects, 3D Gaussian Splatting (3D-GS) [17] has emerged as an efficient representation that can be rendered at high speed on the GPU.
+It has departed from the implicit scene representation using MLPs and instead opted for an explicit representation using Gaussian functions.
+As a result, 3D-GS avoids the expensive cost of densely sampling points in volume rendering, thus achieving real-time performance.
+Given a set of images with the camera parameters calibrated by SfM [33], 3D-GS aims to optimize a set of 3D Gaussians as graphical primitives to explicitly represent the scene.
+To render 2D images, these 3D Gaussians are projected onto the image plane ($z=1$ plane) via a projection function denoted as $\phi$ for differentiable rasterization.
+Unfortunately, Gaussian functions are closed under affine transformations but not under projective transformations.
+Therefore, 3D-GS adopts a local affine approximation [46], specifically approximating the projection function with the first two terms of its Taylor expansion.
+Nevertheless, approximations introduce errors, and the local affine approximation similarly contributes to these errors which may lead to artifacts in the rendered images.
+This paper exploits the relationship between the error of 3D-GS and the Gaussian mean through the analysis of the Taylor remainder term.
 Additionally, it identifies the circumstances under which the error is minimized by determining the extremum of the error function.
-Finally, based on the extremum analysis of the error function, we propose an optimal projection method.
+Finally, based on the extremum analysis of the error function, we propose a novel optimal projection method to minimize the projection error in 3D-GS.
+This method requires only minor code modifications, does not affect real-time rendering performance, and yet achieves a significant improvement in rendering quality.
 Specifically, we project along the direction from the Gaussian mean to the camera center, where the projection plane is tangent to the line connecting the Gaussian mean and the camera center.
-Further validation has been conducted on diverse datasets, comparing the proposed projection with the straightforward approach of projecting 3D Gaussians onto the $z=1$ plane.
-The results illustrate a reduced expected error, leading to real-time, realistic rendering, as shown in Figure 5 and Table 1.
+Compared with the straightforward approach of projecting 3D Gaussians onto the image plane of the camera, the proposed projection reduces expected error while maintaining real-time and realistic rendering performance.
+Furthermore, we find that the projection errors in 3D-GS tend to escalate with the continuous expansion of the field of view and the corresponding reduction in focal length.
+This increasing error leads to more and larger elongated Gaussians, as well as cloud-like Gaussian artifacts, consequently degrading the overall image quality.
+In contrast, our method exhibits greater realism and robustness.
+As illustrated in Figure 1, our method significantly reduces blurriness and artifacts caused by the projection error.
+Through experiments on various datasets and scenes, our method consistently outperforms the original projection method, particularly in settings with short focal lengths.
+And the method is easily implemented and adaptable to various camera models.
 To summarize, we provide the following contributions:
-  • We analyze error of 3D Gaussian Splatting that may lead to artifacts and consequently degrades rendering quality. We have identified the correlation between this error and the Gaussian position.
+  • We provide thorough error analysis for 3D-GS that may lead to artifacts and consequently degrade rendering quality. We have identified the correlation between this error and the Gaussian position.
   • We derive the mathematical expectation of this error function and analyze when this function takes extrema through methods of function optimization.
-  • Based on the conclusion regarding the extremum of the expectation of this error function, we propose a new optimal projection method.
-Instead of naively projecting all Gaussians onto the same plane, we now adopt different tangent plane projections based on the directions from each Gaussian mean to the camera center.
-Through experiments, we further validated the superiority of this method.
-In various datasets and scenes, it consistently achieves better overall performance compared to the original projection method.
+  • We propose a novel optimal projection capable of adapting to diverse camera models that adopts different tangent plane projections based on the directions from each Gaussian mean to the camera center instead of naively projecting all Gaussians onto the same plane
 2 Related Work
-To begin with, we will provide a succinct overview of traditional approaches in novel view synthesis and scene reconstruction.
-Subsequently, we will delve into works related to implicit scene representation using Neural Radiance Fields and explicit scene representation through 3D Gaussian Splatting.
-2.1 Traditional Scene Reconstruction and Rendering
-Early novel-view synthesis methods were based on light fields  [10, 19], transitioning from densely sampled to unstructured capture  [3].
-Structure-from-Motion (SfM) [33] introduced a new era using photo collections for view synthesis.
-Multi-view stereo (MVS) and subsequent algorithms  [9, 5] blended input images but faced challenges.
-Recent neural rendering algorithms [18] have overcome these issues, outperforming traditional methods.
+Novel view synthesis is a longstanding challenge in computer vision and graphics.
+From traditional techniques  [33, 12, 20, 3, 35, 11, 6], to neural network-based scene representations [16, 31, 10, 26, 30, 34, 19], various approaches have struggled to address the problem of synthesizing a new view from captured images.
 2.1 Neural Radiance Fields
@@
 These enhancements enable the handling of multi-resolution image inputs, multi-illumination with occlusion image inputs, sparse-view inputs, and achieve nearly real-time rendering capabilities, respectively.
+Barron et al.
+introduced MipNeRF360 [2] as an extension of MipNeRF [1] to address the issue of generating low-quality renderings for unbounded scenes in NeRF.
 Importantly, these methods thoroughly exploit the intrinsic capabilities of NeRF as an implicit scene representation without introducing additional model priors.
-However, these methods are constrained by the dense sampling of rays, preventing them from achieving real-time performance.
-In contrast, the recent 3D Gaussian Splatting [16] achieves real-time rendering by breaking away from MLPs and ray sampling constraints.
+However, due to the implicit representation of scenes and the dense sampling of points along rays, they still face challenges in achieving real-time performance.
 2.2 3D Gaussian Splatting
-3D Gaussian Splatting [16] utilizes Structure-from-Motion (SFM) [31] point clouds as input to construct an explicit scene representation using Gaussian functions.
-Subsequently, these Gaussian functions undergo a series of coordinate transformations and are projected onto the image plane at $z=1$, where rasterization takes place.
-The significant breakthrough in performance is achieved by eschewing MLPs and ray sampling in favor of Gaussian functions and Gaussian Splatting.
-This shift has garnered widespread attention in the industry, leading to various works building upon this method.
-Gaussian Splatting is applied in SLAM (Simultaneous Localization and Mapping) systems [23].
-Additionally, there is scene editing [6] and segmentation [4] based on 3D Gaussian.
-There are also efforts to integrate 3D Gaussian Splatting with the popular diffusion models [13, 34, 35] in the AIGC domain [20].
-The diffusion model is employed to provide geometric priors for 3D Gaussian Splatting, correcting inconsistencies in multi-view geometry.
-Another method involves tracking dynamic Gaussians through continuous dynamic view synthesis [21].
-However, these enhancement efforts have relatively less emphasis on the 3D Gaussian Splatting method itself and often involve the use of additional models or priors to assist 3D-GS.
-This paper aims to explore the potential of 3D Gaussian Splatting by analyzing the errors that may arise during the projection process.
+3D Gaussian Splatting (3D-GS) [17] has demonstrated notable advancements in rendering performance by departing from MLPs and ray sampling, opting instead for Gaussian functions and Gaussian Splatting.
+This paradigm shift has attracted considerable attention within the industry, leading to various studies building upon this methodology [7].
+3D-GS finds application in SLAM systems [25], scene editing [8] and segmentation [4].
+Efforts [21] have also been made to integrate 3D-GS with popular diffusion models [15, 36, 37] and dynamic scenes [23].
+Recent endeavors aim to enhance the robustness in sparse-view scenarios [40, 42], performance, storage efficiency [29] and mesh reconstruction [13] of 3D-GS.
+However, these improvements do not specifically address errors associated with Gaussian projection.
+Potential methods for enhancing Gaussian functions include [44, 32].
+Nevertheless, they introduce a Gaussian filter and a novel Gaussian depth sorting algorithm, respectively, without investigating the projection function and associated errors.
+These approaches still rely on projecting onto the $z=1$ plane.
+This paper aims to explore the potential of 3D Gaussian Splatting by analyzing errors that may arise during the projection process.
 The analysis will delve into the factors contributing to these errors and propose methods to mitigate and reduce them.
 3 Preliminaries
-(a) 0.095
-(b) 0.95
-Figure 2: Visualizations of the 3D Gaussian Splatting error function $\epsilon\left(\theta_{\mu},\phi_{\mu}\right)$.
-The two subplots from left to right respectively depict the function graphs within the domains of $\left\{\left(\theta_{\mu},\phi_{\mu}\right)\mid\theta\in\left[-\pi/4\cdot 0.095,\pi/4\cdot 0.095\right]~\land~\phi\in\left[-\pi/4\cdot 0.095,\pi/4\cdot 0.095\right]\right\}$ and $\left\{\left(\theta_{\mu},\phi_{\mu}\right)\mid\theta\in\left[-\pi/4\cdot 0.95,\pi/4\cdot 0.95\right]~\land~\phi\in\left[-\pi/4\cdot 0.95,\pi/4\cdot 0.95\right]\right\}$ where $\pi/4$ corresponds to the integration region of the mathematical expectation.
-These representations are employed for a better understanding of the nature of this error function.
-In the world coordinate system, let there be a 3D Gaussian function $G$, characterized by its mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$, expressed as:
-$G_{\mathbf{\mu},\boldsymbol{\Sigma}}\left(\mathbf{x}\right)=\exp\left(-1/2\left(\mathbf{x}-\boldsymbol{\mu}\right)^{\top}\boldsymbol{\Sigma}^{-1}\left(\mathbf{x}-\boldsymbol{\mu}\right)\right)\text{.}$ (1)
-This Gaussian is multiplied by $\alpha$ in blending process.
-To project the Gaussian onto the image plane, the initial step in 3D Gaussian Splatting involves an affine transformation of this Gaussian function from the world coordinate system to the camera coordinate system via a viewing transformation matrix $\mathbf{W}$.
-The transformed Gaussian function is as follows:
-$\begin{split}G\left(\mathbf{x}\right)&=\exp\left(-1/2\left(\mathbf{x}-\boldsymbol{\mu}\right)^{\top}\boldsymbol{\Sigma}^{-1}\left(\mathbf{x}-\boldsymbol{\mu}\right)\right)\\ &=\exp\left(-1/2\left(\mathbf{W}\mathbf{x}-\mathbf{W}\boldsymbol{\mu}\right)^{\top}\left(\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}\right)^{-1}\left(\mathbf{W}\mathbf{x}-\mathbf{W}\boldsymbol{\mu}\right)\right)\end{split}$ (2)
-which can also be represented by the following function:
-$\begin{split}G^{{}^{\prime}}\left(\mathbf{x^{{}^{\prime}}}\right)&=\exp\left(-1/2\left(\mathbf{x^{{}^{\prime}}}-\boldsymbol{\mu^{{}^{\prime}}}\right)^{\top}{\boldsymbol{\Sigma}^{{}^{\prime}}}^{-1}\left(\mathbf{x^{{}^{\prime}}}-\boldsymbol{\mu^{{}^{\prime}}}\right)\right)\end{split}$ (3)
-where $G^{{}^{\prime}}$, $\mathbf{x}^{{}^{\prime}}=\mathbf{W}\mathbf{x}$, $\boldsymbol{\mu}^{{}^{\prime}}=\mathbf{W}\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}^{{}^{\prime}}=\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}$ represent the 3D Gaussian function, point, mean and covariance matrix in the camera coordinate system, respectively.
-Subsequently, it is necessary to project the Gaussian function in the camera coordinate system onto the $z=1$ plane. The projection plane is formulated as:
-$\mathbf{x_{0}}^{\top}\cdot\left(\mathbf{x}^{{}^{\prime}}-\mathbf{x_{0}}\right)=0$ (4)
-where $\mathbf{x_{0}}=\left[\begin{matrix}0,0,1\end{matrix}\right]^{\top}$ represents the projection of the camera coordinate system origin onto this plane.
-According to the projection plane equation (Equation 4), the projection function $\phi$ is obtained as:
-$\phi\left(\mathbf{x}^{{}^{\prime}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{0}}^{T}\mathbf{x}^{{}^{\prime}}\right)^{-1}\left(\mathbf{x_{0}}^{T}\mathbf{x_{0}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{0}}^{T}\mathbf{x}^{{}^{\prime}}\right)^{-1}.$ (5)
+Similar to NeRF [27], the input to 3D-GS consists of a set of images, together with the corresponding cameras calibrated by SFM [33].
+However, in contrast to NeRF, 3D-GS takes the sparse point cloud generated during the SFM calibration process as input.
+From these points, it constructs a set of 3D Gaussians as graphical primitives explicitly representing the scene.
+Each 3D Gaussian is characterized by its position (mean) $\boldsymbol{\mu}$, covariance matrix $\boldsymbol{\Sigma}$, and opacity $\alpha$, and carries anisotropic spherical harmonics (SH) representing directional appearance component (color) of the radiance field, following standard practice [28, 9].
+Subsequently, these 3D Gaussians are projected onto the image plane ($z=1$ plane) via a projection function for differentiable rasterization.
+Next, we will provide a detailed explanation of the Gaussian projection process in 3D-GS.
+In the world space, we define a 3D Gaussian function by $G$, characterized by its mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$.
+To project the Gaussian onto the image plane, the initial step in 3D-GS involves an affine transformation of this Gaussian function from the world space to the camera space via a viewing transformation matrix $\mathbf{W}$, yielding a transformed Gaussian function:
+$\begin{split}G\left(\mathbf{x}\right)=\exp\left\{-1/2\left(\mathbf{W}\mathbf{x}-\mathbf{W}\boldsymbol{\mu}\right)^{\top}\left(\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}\right)^{-1}\left(\mathbf{W}\mathbf{x}-\mathbf{W}\boldsymbol{\mu}\right)\right\}\text{.}\end{split}$ (1)
+We denote $G^{{}^{\prime}}$ as the 3D Gaussian in the camera space, with $\mathbf{x}^{{}^{\prime}}=\mathbf{W}\mathbf{x}$, $\boldsymbol{\mu}^{{}^{\prime}}=\mathbf{W}\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}^{{}^{\prime}}=\mathbf{W}\boldsymbol{\Sigma}\mathbf{W}^{\top}$ as the point, mean and covariance matrix in the camera space, respectively.
+Subsequently, traditional 3D-GS requires projecting the Gaussian function in the camera space onto the $z=1$ plane, with the projection function $\varphi$ defined as:
+$\varphi\left(\mathbf{x}^{{}^{\prime}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{0}}^{\top}\mathbf{x}^{{}^{\prime}}\right)^{-1}\left(\mathbf{x_{0}}^{\top}\mathbf{x_{0}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{0}}^{\top}\mathbf{x}^{{}^{\prime}}\right)^{-1}$ (2)
+where $\mathbf{x_{0}}=\left[\begin{matrix}0,0,1\end{matrix}\right]^{\top}$ represents the projection of the camera space’s origin onto this plane.
 Expanding this projection function to the first order using Taylor series yields:
-$\begin{split}\phi\left(\mathbf{x}^{{}^{\prime}}\right)&=\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)+\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)+R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)\\ &\approx\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)+\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)\end{split}$ (6)
-where $\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$ is the Jacobian of the affine approximation of the projective transformation, denoted as $\mathbf{J}$.
-Applying this local affine approximation by neglecting the Taylor remainder term allows to derive the 2D Gaussian function $G_{\text{2D}}$ projected onto the $z=1$ plane similar to Equation 2:
-$\begin{split}G_{\text{2D}}\left(\mathbf{x}^{{}^{\prime}}\right)&=\exp\left(-1/2\left(\mathbf{J}\mathbf{x}^{{}^{\prime}}-\mathbf{J}\boldsymbol{\mu}^{{}^{\prime}}\right)^{\top}\left(\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}\right)^{-1}\left(\mathbf{J}\mathbf{x}^{{}^{\prime}}-\mathbf{J}\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\\ &\approx\exp\left(-1/2\left(\phi\left(\mathbf{x}^{{}^{\prime}}\right)-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)^{\top}\left(\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}\right)^{-1}\left(\phi\left(\mathbf{x}^{{}^{\prime}}\right)-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\right).\end{split}$ (7)
-Note that since the rank of the matrix $\mathbf{J}$ is 2, the inverse of $\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}$ is, in fact, the inverse of the covariance matrix of the 2D Gaussian which means skipping the third row and column of this matrix.
-Similarly, this function can be expressed in another form:
-$G_{\text{2D}}\left(\mathbf{x_{\text{2D}}}\right)=\exp\left(-1/2\left(\mathbf{x_{\text{2D}}}-\boldsymbol{\mu_{\text{2D}}}\right)^{\top}\boldsymbol{\Sigma}_{\text{2D}}^{-1}\left(\mathbf{x_{\text{2D}}}-\boldsymbol{\mu_{\text{2D}}}\right)\right)$ (8)
-where $G_{\text{2D}}$, $\mathbf{x}_{\text{2D}}=\phi\left(\mathbf{x}^{{}^{\prime}}\right)$, $\boldsymbol{\mu}_{\text{2D}}=\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$ and $\boldsymbol{\Sigma}_{\text{2D}}=\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}$ represent the 2D Gaussian function, point, mean and $\times 2$ 2 covariance matrix in the image coordinate system, respectively.
-After the aforementioned steps, 3D Gaussian Splatting projects the 3D Gaussian from the world coordinate system onto the image plane.
-Subsequently, rasterization is performed on the image plane to obtain the rendered image.
+$\begin{split}\varphi\left(\mathbf{x}^{{}^{\prime}}\right)&=\varphi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)+\frac{\partial\varphi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)+R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)\\ &\approx\varphi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)+\frac{\partial\varphi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)\end{split}$ (3)
+where $\frac{\partial\varphi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$ is the Jacobian of the affine approximation of the projective transformation, denoted as $\mathbf{J}$.
+Applying this local affine approximation by neglecting the Taylor remainder term allows to derive the 2D Gaussian function $G_{\text{2D}}$ projected onto the $z=1$ plane as:
+$\begin{split}G_{\text{2D}}\left(\mathbf{x}^{{}^{\prime}}\right)&=\exp\left\{-1/2\left(\mathbf{J}\mathbf{x}^{{}^{\prime}}-\mathbf{J}\boldsymbol{\mu}^{{}^{\prime}}\right)^{\top}\left(\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}\right)^{-1}\left(\mathbf{J}\mathbf{x}^{{}^{\prime}}-\mathbf{J}\boldsymbol{\mu}^{{}^{\prime}}\right)\right\}\\ &\approx\exp\left\{-1/2\left(\varphi\left(\mathbf{x}^{{}^{\prime}}\right)-\varphi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)^{\top}\left(\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}\right)^{-1}\left(\varphi\left(\mathbf{x}^{{}^{\prime}}\right)-\varphi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\right\}.\end{split}$ (4)
+Since the rank of the matrix $\mathbf{J}$ is 2, the inverse of $\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}$ is, in fact, the inverse of the covariance matrix of the 2D Gaussian obtained by skipping the third row and column [46].
+We denote $\mathbf{x}_{\text{2D}}=\varphi\left(\mathbf{x}^{{}^{\prime}}\right)$, $\boldsymbol{\mu}_{\text{2D}}=\varphi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$ and $\boldsymbol{\Sigma}_{\text{2D}}=\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}$ as the point, mean and $2\times 2$ covariance matrix of $G_{\text{2D}}$ in the image space, respectively.
+For a pixel $(u,v)$, we obtain the value of the 2D Gaussian function by querying $\mathbf{x_{\text{2D}}}=\left[\begin{matrix}u,v\end{matrix}\right]^{\top}$, and then perform alpha blending to derive the color of the pixel.
+The blue box in Figure 3 illustrates the projection process.
 4 Local Affine Approximation Error
@@
 the local affine approximation.
-The rasterization uses a 2D Gaussian function that is not the true projection function of the 3D Gaussian.
+Consequently, 3D-GS actually adopts a 2D Gaussian function that is not the true projection of the initial 3D Gaussian.
 The error introduced by this approximation can be characterized by the Taylor remainder term in Equation 3:
-$\begin{split}&R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)=\phi\left(\mathbf{x}^{{}^{\prime}}\right)-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)-\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)\end{split}$ (9)
+$\begin{split}&R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)=\varphi\left(\mathbf{x}^{{}^{\prime}}\right)-\varphi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)-\frac{\partial\varphi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\left(\mathbf{x}^{{}^{\prime}}-\boldsymbol{\mu}^{{}^{\prime}}\right)\end{split}$ (5)
 where
-$\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\mathbb{I}\otimes\left(\mathbf{x_{0}}^{T}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}-\mathbf{x_{0}}\left(\mathbf{x_{0}}^{T}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{T}\mathbf{x_{0}}\right)^{-1}{\boldsymbol{\mu}^{{}^{\prime}}}^{T}$ (10)
-and $\mathbb{I}$ represents the identity matrix; $\otimes$ denotes the multiplication of a matrix and a scalar.
-It is not difficult to observe that this Taylor remainder term is a three-dimensional vector related to the random variable $\mathbf{x}^{{}^{\prime}}$ and the mean $\boldsymbol{\mu}^{{}^{\prime}}$ of the Gaussian function.
-Therefore, by computing the square of the Frobenius norm of this vector $\|R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)\|_{F}^{2}$ and taking the mathematical expectation of this norm function with respect to the random variable $\mathbf{x}^{{}^{\prime}}$, we ultimately obtain an error function that depends solely on the Gaussian mean $\boldsymbol{\mu}^{{}^{\prime}}$:
+$\frac{\partial\varphi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\mathbb{I}\otimes\left(\mathbf{x_{0}}^{\top}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}-{\boldsymbol{\mu}^{{}^{\prime}}}\left(\mathbf{x_{0}}^{\top}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{\top}\mathbf{x_{0}}\right)^{-1}\mathbf{x_{0}}^{\top}\text{.}$ (6)
+$\mathbb{I}$ represents the identity matrix and $\otimes$ denotes the multiplication of a matrix and a scalar.
+Clearly, this Taylor remainder term is a three-dimensional vector related to the random variable $\mathbf{x}^{{}^{\prime}}$ and the mean $\boldsymbol{\mu}^{{}^{\prime}}$ of the Gaussian function.
+Therefore, by computing the square of the Frobenius norm of this vector $\|R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)\|_{F}^{2}$ and taking the mathematical expectation of this norm function with respect to the random variable $\mathbf{x}^{{}^{\prime}}$, we obtain an error function that depends solely on the Gaussian mean $\boldsymbol{\mu}^{{}^{\prime}}$:
 $\epsilon\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\int\limits_{\mathbf{x}^{{}^{\prime}}\in\mathcal{X}^{{}^{\prime}}}\|R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)\|_{F}^{2}\mathrm{d}\mathbf{x}^{{}^{\prime}}\text{.}$ (7)
-Before seeking the extremum of the error function, let’s simplify the integral expression.
-First, we simplify $\mathbf{x}^{{}^{\prime}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$ to unit vectors projected onto the unit sphere centered at the camera center (hereafter referred to as the unit sphere) by proving that the composition of transformations $\left(\phi\circ\pi\right)$, projecting a point in three-dimensional space onto the unit sphere $\pi$ and subsequently onto the projection plane $\phi$, is equivalent to directly projecting a point in three-dimensional space onto the projection plane.
-This simplification ensures that the entire error function involves only operations between three unit space vectors.
-The proof proceeds as follows:
-$\begin{split}\left(\phi\circ\pi\right)\left({\mathbf{x}^{{}^{\prime}}}\right)&=\phi\left({\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\right)\\ &={\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\left(\mathbf{x_{0}}^{\top}\left({\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\right)\right)^{-1}\\ &={\mathbf{x}^{{}^{\prime}}}\left(\mathbf{x_{0}}^{T}{\mathbf{x}^{{}^{\prime}}}\right)^{-1}\\ &=\phi\left({\mathbf{x}^{{}^{\prime}}}\right)\text{.}\end{split}$ (12)
-The simplified error function $\epsilon$ involves three unit vectors, or equivalently, three points on the unit sphere: $\mathbf{x}^{{}^{\prime}}$, $\mathbf{x_{0}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$, as illustrated in Figure 3.
+The simplified error function $\epsilon$ involves three unit vectors: $\mathbf{x}^{{}^{\prime}}$, $\mathbf{x_{0}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$.
 Their spherical coordinates are given as follows:
 $\mathbf{x_{0}}=\left[\begin{matrix}\sin{\left(\phi_{0}\right)}\cos{\left(\theta_{0}\right)}\\ -\sin{\left(\theta_{0}\right)}\\ \cos{\left(\phi_{0}\right)}\cos{\left(\theta_{0}\right)}\end{matrix}\right],\mathbf{x}^{{}^{\prime}}=\left[\begin{matrix}\sin{\left(\phi\right)}\cos{\left(\theta\right)}\\ -\sin{\left(\theta\right)}\\ \cos{\left(\phi\right)}\cos{\left(\theta\right)}\end{matrix}\right],\boldsymbol{\mu}^{{}^{\prime}}=\left[\begin{matrix}\sin{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}\\ -\sin{\left(\theta_{\mu}\right)}\\ \cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}\end{matrix}\right]$ (8)
-where $\phi_{0}=0$ and $\theta_{0}=0$. Based on Equation 9, Equation 10, and Equation 13, the following derivations can be obtained:
-$\mathbf{J}=\frac{\partial\phi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\left[\begin{matrix}\frac{1}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}&0&-\frac{\sin{\left(\phi_{\mu}\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}\\ 0&\frac{1}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}&\frac{\sin{\left(\theta_{\mu}\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\theta_{\mu}\right)}}\\ 0&0&0\end{matrix}\right]$ (14)
-$R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)=\left[\begin{matrix}-\frac{\sin{\left(\phi-\phi_{\mu}\right)}\cos{\left(\theta\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}+\tan{\left(\phi\right)}-\tan{\left(\phi_{\mu}\right)}\\ \frac{\sin{\left(\theta\right)}}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}-\frac{\sin{\left(\theta_{\mu}\right)}\cos{\left(\phi\right)}\cos{\left(\theta\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\theta_{\mu}\right)}}+\frac{\tan{\left(\theta_{\mu}\right)}}{\cos{\left(\phi_{\mu}\right)}}-\frac{\tan{\left(\theta\right)}}{\cos{\left(\phi\right)}}\\ 0\end{matrix}\right]\text{.}$ (15)
+where $\phi_{0}=0$ and $\theta_{0}=0$. Please refer to the detailed proofs in the supplementary materials.
 The size of the integration region is related to the covariance of the Gaussian function.
 Since the current discussion specifically focuses on the impact of the mean on the error rather than the variance, let’s assume the integration region to be $\left\{\mathbf{x}^{{}^{\prime}}\mid\theta\in\left[-\pi/4+\theta_{\mu},\pi/4+\theta_{\mu}\right]~{}\land~{}\phi\in\left[-\pi/4+\phi_{\mu},\pi/4+\phi_{\mu}\right]\right\}$.
-Then, substituting Equation 15 into Equation 11, we can derive:
+Then, based on Equation 5-8, we obtain:
 $\begin{split}\epsilon\left(\theta_{\mu},\phi_{\mu}\right)&=\int_{-\pi/4+\theta_{\mu}}^{\pi/4+\theta_{\mu}}\int_{-\pi/4+\phi_{\mu}}^{\pi/4+\phi_{\mu}}\left(-\frac{\sin{\left(\phi-\phi_{\mu}\right)}\cos{\left(\theta\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}+\tan{\left(\phi\right)}-\tan{\left(\phi_{\mu}\right)}\right)^{2}+\\ &\left(\frac{\sin{\left(\theta\right)}}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}-\frac{\sin{\left(\theta_{\mu}\right)}\cos{\left(\phi\right)}\cos{\left(\theta\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\theta_{\mu}\right)}}+\frac{\tan{\left(\theta_{\mu}\right)}}{\cos{\left(\phi_{\mu}\right)}}-\frac{\tan{\left(\theta\right)}}{\cos{\left(\phi\right)}}\right)^{2}\mathrm{d}\theta\mathrm{d}\phi\text{.}\end{split}$ (9)
-The function is partially differentiated with respect to $\theta_{\mu}$ and $\phi_{\mu}$. It is observed that at $\theta_{\mu}=\theta_{0}=0$ and $\phi_{\mu}=\phi_{0}=0$, the following holds:
+Fortunately, this integral has a close-form expression, but its form is very complicated. We report the close-form expression in the supplemental materials and visually depict it in Figure 2.
+Figure 2: Visualization of the 3D Gaussian Splatting error function $\epsilon\left(\theta_{\mu},\phi_{\mu}\right)$ under two distinct domains.
+The domain of the function is $\left\{\left(\theta_{\mu},\phi_{\mu}\right)\mid\theta_{\mu}\in\left[-\lambda\pi/4,\lambda\pi/4\right]~{}\land~{}\phi_{\mu}\in\left[-\lambda\pi/4,\lambda\pi/4\right]\right\}$ where $\left(\theta_{\mu},\phi_{\mu}\right)$ represents the polar coordinate of the 3D Gaussian mean and $\lambda$ represents the scaling factor of the domain.
+The error function is partially differentiated with respect to $\theta_{\mu}$ and $\phi_{\mu}$.
+It is easy to check that at $\theta_{\mu}=\theta_{0}=0$ and $\phi_{\mu}=\phi_{0}=0$, the following holds:
 $\frac{\partial\epsilon}{\partial\theta_{\mu}}\left(0,0\right)=0,\frac{\partial\epsilon}{\partial\phi_{\mu}}\left(0,0\right)=0\text{.}$ (10)
-Therefore, this point is an extremum point of the function.
-Upon verification, it is confirmed that this point is the minimum value point of the function.
-We have visualized the error function, as shown in Figure 2.
-Through Subfigure 2a in Figure 2, it can be observed that this function is a concave function, and it attains its minimum value at (0,0)(0,0)( 0 , 0 ).
-Examining Subfigure 2b in Figure 2, it is evident that in the majority of the region close to the origin, the function values are not significantly different.
-However, as the function approaches the integration limits, the function values increase rapidly, resulting in a substantial difference between the maximum and minimum values.
-In other words, in most cases, the error is small and does not significantly affect the quality of rendered images.
-This is why 3D Gaussian Splatting utilizes local affine approximation but still successfully reconstructs the scene to obtain a high-quality novel view image.
-Nevertheless, the naive projection of all gaussians onto the same plane $z=1$ in 3D Gaussian Splatting may lead to larger projection errors for gaussians farther from the plane center (the projection point of the camera center to the plane), causing artifacts, as shown in Figure 5.
-Figure 3: Toy 2D visualizations of the projection transformation.
-$\mathbf{x}^{{}^{\prime}}$, $\mathbf{x_{0}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$ are all simplified unit vectors, representing the random variable, the tangency point of the projection plane and the unit sphere, and the Gaussian mean, respectively, denoted by green, purple, and blue arrows in the diagram.
-Due to the $3\sigma$ rule in the Gaussian distribution and the fact that actual rasterization does not sample all points, we represent the integration domain with a circular arc symmetric about the mean vector $\boldsymbol{\mu}^{{}^{\prime}}$.
+Therefore, this point is actually an extremum point of the function.
+We also confirm that this point is the minimum of the function.
+Indeed, from Figure 2(a), it is also apparent that the function is concave, with its minimum value occurring at $(0,0)$.
+Additionally, it can be observed that the minimum value of the error function is greater than zero, due to the fact that Gaussian functions are not closed under projection transformations.
+Figure 2(b) shows that in the majority of the region close to the origin, the function values vary very gently.
+However, as the independent variable of the function approaches the boundaries, the function values increase rapidly, resulting in a substantial difference between the maximum and minimum values.
+The above analysis indicates that the error is small and does not significantly affect the quality of rendered images.
+Therefore, traditional 3D-GS [17] utilizes local affine approximation but still successfully reconstructs the scene to obtain a high-quality novel view image.
+Nevertheless, the naive projection of all Gaussians onto the same plane $z=1$ in 3D-GS may lead to larger projection errors for Gaussians farther from the plane center (the projection point of the camera center to the plane), causing artifacts.
+These artifacts are severe when using wide-angle lenses, significantly degrading the quality of the renderings.
 5 Optimal Gaussian Splatting
-Figure 4: Illustration of the rendering pipeline for Optimal Gaussian Splatting.
+Figure 3: Illustration of the rendering pipeline for our Optimal Gaussian Splatting and the projection of 3D-GS [17].
+The blue box depicts the projection process of the original 3D-GS, which straightforwardly projects all Gaussians onto the same projection plane.
+In contrast, the red box illustrates our approach, where we project individual Gaussians onto corresponding tangent planes.
+Based on the analysis of the error function, we introduce Optimal Gaussian Splatting shown in Figure 3.
+This helps us to achieve smaller projection errors, resulting in higher-quality rendering.
 The first step, similar to the original 3D-GS [17], involves transforming world coordinates into the camera coordinate system.
@@
 Finally, for each pixel in the image, rays are cast onto the unit sphere to retrieve the color of that pixel, resulting in the rendered image.
-Based on the analysis of the error function, in this section, we will introduce Optimal Gaussian Splatting shown in Figure 4 that exhibits smaller projection errors, resulting in higher-quality rendering, as shown in Figure 5 and Table 1.
 5.1 Optimal Projection
-Through the analysis of the error function in Section 4, we discovered that the error function attains its minimum value when the projection of the Gaussian mean on the plane coincides with the projection from the camera center to the plane.
+We discover that the error function attains its minimum value when the projection of the Gaussian mean on the plane coincides with the projection from the camera center to the plane.
 Therefore, in Optimal Gaussian Splatting, we employ an optimal projection method.
@@
 These projection planes are determined by the tangent planes formed by the Gaussian mean and the line connecting it to the camera center.
-The specific projection plane is ormulated as:
-$\mathbf{x^{\mu}}^{\top}\cdot\left(\mathbf{x}^{{}^{\prime}}-\mathbf{x^{\mu}}\right)=0$ (18)
+The specific projection plane is formulated as:
+$\mathbf{x_{p}}^{\top}\cdot\left(\mathbf{x}^{{}^{\prime}}-\mathbf{x_{p}}\right)=0$ (11)
 where
-$\mathbf{x^{\mu}}=\pi\left({\boldsymbol{\mu}^{{}^{\prime}}}\right)={\boldsymbol{\mu}^{{}^{\prime}}}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{\top}{\boldsymbol{\mu}^{{}^{\prime}}}\right)^{-1/2}$ (19)
-represents the projection of the camera coordinate system origin onto this plane.
-According to the projection plane equation (Equation 18 and Equation19), the optimal projection function $\phi_{o}$ is obtained as:
-$\phi^{\mu}\left(\mathbf{x}^{{}^{\prime}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x^{\mu}}^{T}\mathbf{x}^{{}^{\prime}}\right)^{-1}\left(\mathbf{x^{\mu}}^{T}\mathbf{x^{\mu}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x^{\mu}}^{T}\mathbf{x}^{{}^{\prime}}\right)^{-1}\text{.}$ (20)
-The corresponding local affine approximation Jacobian matrix $\mathbf{J_{o}}$ is:
-$\mathbf{J^{\mu}}=\frac{\partial\phi^{\mu}}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\mathbb{I}\otimes\left(\mathbf{x^{\mu}}^{T}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}-\mathbf{x^{\mu}}\left(\mathbf{x^{\mu}}^{T}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{T}\mathbf{x^{\mu}}\right)^{-1}{\boldsymbol{\mu}^{{}^{\prime}}}^{T}\text{.}$ (21)
+$\mathbf{x_{p}}=\varpi\left({\boldsymbol{\mu}^{{}^{\prime}}}\right)={\boldsymbol{\mu}^{{}^{\prime}}}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{\top}{\boldsymbol{\mu}^{{}^{\prime}}}\right)^{-1/2}$ (12)
+represents the projection of the camera space’s origin onto this plane and $\varpi$ function represents the transformation that projects points onto the unit sphere.
+According to Equation 11, the optimal projection function $\varphi_{\text{p}}$ is obtained as:
+$\varphi_{\text{p}}\left(\mathbf{x}^{{}^{\prime}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{p}}^{\top}\mathbf{x}^{{}^{\prime}}\right)^{-1}\left(\mathbf{x_{p}}^{\top}\mathbf{x_{p}}\right)=\mathbf{x}^{{}^{\prime}}\left(\mathbf{x_{p}}^{\top}\mathbf{x}^{{}^{\prime}}\right)^{-1}\text{.}$ (13)
+The corresponding local affine approximation Jacobian matrix $\mathbf{J_{p}}$ is:
+$\mathbf{J_{p}}=\frac{\partial\varphi_{\text{p}}}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\mathbb{I}\otimes\left(\mathbf{x_{p}}^{\top}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}-{\boldsymbol{\mu}^{{}^{\prime}}}\left(\mathbf{x_{p}}^{\top}\boldsymbol{\mu}^{{}^{\prime}}\right)^{-1}\left({\boldsymbol{\mu}^{{}^{\prime}}}^{\top}\mathbf{x_{p}}\right)^{-1}\mathbf{x_{p}}^{\top}\text{.}$ (14)
 When $\boldsymbol{\mu}^{{}^{\prime}}=\left[\begin{matrix}\mu_{x},\mu_{y},\mu_{z}\end{matrix}\right]^{\top}$, the specific form of this Jacobian matrix is:
-$\mathbf{J^{\mu}}=\left[\begin{matrix}\frac{\mu_{y}^{2}+\mu_{z}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{x}\mu_{y}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{x}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\\ -\frac{\mu_{x}\mu_{y}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&\frac{\mu_{x}^{2}+\mu_{z}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{y}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\\ -\frac{\mu_{x}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{y}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&\frac{\mu_{x}^{2}+\mu_{y}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\end{matrix}\right]$ (22)
+$\mathbf{J_{p}}=\left[\begin{matrix}\frac{\mu_{y}^{2}+\mu_{z}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{x}\mu_{y}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{x}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\\ -\frac{\mu_{x}\mu_{y}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&\frac{\mu_{x}^{2}+\mu_{z}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{y}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\\ -\frac{\mu_{x}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&-\frac{\mu_{y}\mu_{z}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}&\frac{\mu_{x}^{2}+\mu_{y}^{2}}{\left(\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}\right)^{\frac{3}{2}}}\end{matrix}\right]\text{.}$ (15)
+It’s very easy to implement modifications to the Jacobian matrix in the forward process.
 5.2 Unit Sphere Based Rasterizer
-For Optimal Projection, we have employed a Unit Sphere Based Rasterizer for rasterization to generate images.
+For the proposed Optimal Projection, we have proposed a Unit Sphere Based Rasterizer to generate images.
 Through Optimal Projection, we obtain the projection of the three-dimensional Gaussians on the tangent plane of the unit sphere instead of obtaining the Gaussians in the image space.
-Therefore, we need to rasterize based on this unit sphere to generate images.
-In specific terms, for a pixel $\left(u,v\right)$ on the image, similar to NeRF, we cast a ray.
+Therefore, we need to rasterize based on this unit sphere.
+Specifically, for a pixel $\left(u,v\right)$ on the image, similar to NeRF, we cast a ray.
 However, unlike NeRF, we do not involve the extensive sampling of points along the ray, which would significantly degrade performance.
 Instead, our focus is solely on determining which tangent-plane Gaussians the ray intersects on the unit sphere:
-$\mathbf{x^{\mu}_{\text{2D}}}=\phi^{\mu}\left(\left[\begin{matrix}\left(u-c_{x}\right)/f_{x}\\ \left(v-c_{y}\right)/f_{y}\\ 1\end{matrix}\right]\right)$ (23)
-We then query the function values of these Gaussians for alpha blending to obtain the color
-$G^{\mu}_{\text{2D}}\left(\mathbf{x^{\mu}_{\text{2D}}}\right)=\exp\left(-1/2\left(\mathbf{x^{\mu}_{\text{2D}}}-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)^{\top}\left(\mathbf{J^{\mu}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J^{\mu}}^{\top}\right)^{-1}\left(\mathbf{x^{\mu}_{\text{2D}}}-\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\right)$ (24)
-$\alpha\left(u,v,\boldsymbol{\mu}\right)=\alpha_{\boldsymbol{\mu}}\cdot G^{\mu}_{\text{2D}}\left(\mathbf{x^{\mu}_{\text{2D}}}\right)$ (25)
-$\mathbf{C}\left(u,v\right)=\sum_{z=z_{\text{near}}}^{z_{\text{far}}}{\text{SH}\left(\boldsymbol{\mu}_{z}\right)\alpha\left(u,v,\boldsymbol{\mu}_{z}\right)\prod_{k=z_{\text{near}}}^{z}\left(1-\alpha\left(u,v,\boldsymbol{\mu}_{k}\right)\right)}$ (26)
-where the summation and product from $z_{\text{near}}$ to $z_{\text{far}}$ represent discretized depths sorted in ascending order from near to far for the depths of 3D Gaussians encountered by the ray, $\alpha_{\boldsymbol{\mu}}$ represents the transparency of the 3D Gaussian with its mean $\boldsymbol{\mu}$, $\boldsymbol{\mu}_{z}$ represents the mean of the Gaussian at depth $z$, and SH represents the spherical harmonics of this Gaussian.
-It is easy to notice that both $\mathbf{J}^{\mu}$ in Equation 21 and $\mathbf{J}$ in Equation 10 are matrices with a rank of 2.
-Therefore, in fact, $\mathbf{J^{\mu}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J^{\mu}}^{\top}$ does not have an inverse.
-We perform a reversible transformation on $\mathbf{J}^{\mu}$ to make the third row zero.
-Consequently, $\mathbf{J^{\mu}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J^{\mu}}^{\top}$ becomes a matrix with both the third row and the third column being zero.
-This allows us to treat it similar to $\mathbf{J}$, skipping the third row and third column, and compute its inverse as if it were a full-rank $\times 2$ 2 matrix.
-Certainly, to ensure the equation still holds, this invertible matrix also needs to be used for left multiplication with $\mathbf{x^{\mu}_{\text{2D}}}$ and $\phi\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$.
+$\mathbf{x}_{\text{2D}}=\varphi_{\text{p}}\left(\left[\begin{matrix}\left(u-c_{x}\right)/f_{x}\\ \left(v-c_{y}\right)/f_{y}\\ 1\end{matrix}\right]\right)$ (16)
+where $c_{x}$, $c_{y}$, $f_{x}$, $f_{y}$ denote the intrinsic parameters of the pinhole camera. We then query the function values of these Gaussians for alpha blending to obtain the color
+$G_{\text{2D}}\left(\mathbf{x_{\text{2D}}}\right)=\exp\left\{-1/2\left(\mathbf{x_{\text{2D}}}-\varphi_{\text{p}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)^{\top}\left(\mathbf{J_{p}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J_{p}}^{\top}\right)^{-1}\left(\mathbf{x_{\text{2D}}}-\varphi_{\text{p}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)\right)\right\}\text{.}$ (17)
+Note that both $\mathbf{J_{p}}$ in Equation 14 and $\mathbf{J}$ in Equation 6 are matrices with a rank of 2.
+Consequently, the inverse of the covariance matrix of the 2D Gaussian is also obtained by skipping the third row and column.
+However, the third row and third column of $\mathbf{J_{p}}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J_{p}}^{\top}$ are not entirely zeros, as opposed to $\mathbf{J}\boldsymbol{\Sigma}^{{}^{\prime}}\mathbf{J}^{\top}$.
+To ensure the equation still holds, an invertible matrix $\mathbf{Q}$ needs to be used for left multiplication with $\mathbf{x_{\text{2D}}}$, $\varphi_{\text{p}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)$ and $\mathbf{J_{p}}$:
+$\mathbf{Q}=\left[\begin{matrix}\frac{\mu_{z}}{\sqrt{\mu_{x}^{2}+\mu_{z}^{2}}}&0&-\frac{\mu_{x}}{\sqrt{\mu_{x}^{2}+\mu_{z}^{2}}}\\ -\frac{\mu_{x}\mu_{y}}{\sqrt{\mu_{x}^{2}+\mu_{z}^{2}}\sqrt{\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}}}&\frac{\sqrt{\mu_{x}^{2}+\mu_{z}^{2}}}{\sqrt{\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}}}&-\frac{\mu_{y}\mu_{z}}{\sqrt{\mu_{x}^{2}+\mu_{z}^{2}}\sqrt{\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}}}\\ \frac{\mu_{x}}{\sqrt{\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}}}&\frac{\mu_{y}}{\sqrt{\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}}}&\frac{\mu_{z}}{\sqrt{\mu_{x}^{2}+\mu_{y}^{2}+\mu_{z}^{2}}}\end{matrix}\right]\text{.}$ (18)
+Since our projection does not rely on the perspective image plane $z=1$ as in the original 3D-GS [17], we can adapt to different camera models such as fisheye cameras and panoramic images by simply modifying the transformation from image space to camera space, as described in Equation 16.
 6 Experiments
-Table 1: Quantitative evaluation of our method compared to previous work, computed over three datasets.
-The Mip-NeRF360 [2] dataset encompasses both Indoor and Outdoor scenarios, while Tanks.
+To validate the effectiveness of the Optimal Gaussian Splatting derived from theoretical error analysis, we conduct a series of experiments, comparing it with the original 3D-GS [17] and some current state-of-the-art methods.
+6.1 Implementation
+We implemented Optimal Gaussian Splatting based on the PyTorch framework in 3D-GS [17] and wrote custom CUDA kernels for rasterization.
+We used the default parameters of 3D-GS to maintain consistency with the original 3D-GS and prevent other factors from introducing interference into the results.
+6.2 Datasets
+We test our algorithm on a total of 13 real scenes which are the same as those used in the original 3D-GS [17].
+In particular, we evaluate our approach on the complete set of scenes featured in Mip-NeRF360 [2], which currently represents the state-of-the-art in NeRF rendering quality.
+Additionally, we test our method on two scenes from the Tanks & Temples dataset [18] and two scenes provided by Hedman et al.
+[14].
+The selected scenes exhibit diverse capture styles, encompassing both confined indoor environments and expansive, unbounded outdoor settings.
+6.3 Results
+Table 1: Quantitative evaluation of our method compared to previous work across three datasets.
+The Mip-NeRF360 [2] dataset includes both indoor and outdoor scenarios.
+Tanks.
 and Deep.
-indicates the average results over the Tank&\&&Templates [17] and Deep Blending [12] datasets, respectively.
-The term Average denotes the mean value of measurements across all scenarios in these datasets.
+indicates the average results over the Tanks&\&&Templates [18] and Deep Blending [14] datasets.
+Average represents the mean value across all scenarios in these datasets.
 Dataset Average Mip-NeRF360 Tanks. and Deep.
@@
 Ours 27.17 0.836 0.210 27.48 0.821 0.209 26.44 0.872 0.214
-Figure 5: We show comparisons of ours to previous methods and the corresponding ground truth images from held-out test views.
-Due to space constraints, we have selected methods such as 3D-GS [16] and Mip-NeRF360 [2], which are comparable in terms of metrics, for qualitative comparisons.
-The scenes are, from the top down: Truck and Train from Tanks&Temples [17]; Playroom, DrJohnson from the Deep Blending dataset [12] and Room, Bonsai, Counter and Stump from Mip-NeRF360 dataset [2]; Non-obvious Differences in quality highlighted by arrows/insets.
-To further validate the effectiveness of the Optimal Gaussian Splatting derived from theoretical analysis, we conducted a series of experiments, comparing it with the original 3D-GS [16] and some current state-of-the-art methods.
-6.1 Implementation
-We implemented Optimal Gaussian Splatting based on the PyTorch framework in 3D-GS [16] and wrote custom CUDA kernels for rasterization.
-We used the default parameters of 3D-GS to maintain consistency with the original 3D-GS and prevent other factors from introducing interference into the results.
-6.2 Datasets
-We tested our algorithm on a total of 13 real scenes which are the same as those used in the original 3D-GS [16] paper.
-In particular, we evaluated our approach on the complete set of scenes featured in Mip-NeRf360 [2], which currently represents the state-of-the-art in NeRF rendering quality.
-Additionally, we tested our method on two scenes from the Tanks & Temples dataset [17] and two scenes provided by Hedman et al.
-[12].
-The selected scenes exhibit diverse capture styles, encompassing both confined indoor environments and expansive, unbounded outdoor settings.
-In other words, our evaluation includes a wide range of scenes to assess the performance of our method comprehensively.
-6.3 Comparisons
+Figure 4: We show comparisons of our method to previous methods and the corresponding ground truth images from held-out test views.
+The scenes are, from the top down: Truck from Tanks&\&&Temples [18]; Playroom from the Deep Blending dataset [14] and Bonsai, Counter from Mip-NeRF360 dataset [2].
+Differences in quality highlighted by arrows/insets.
+Figure 5: We derive the projection error function with respect to image coordinates $(u,v)$ and focal length through the transformation between image and polar coordinates $(\theta_{\mu},\phi_{\mu})$.
+We visualize this function, showing 3D-GS rendered images with long and short focal lengths, followed by the corresponding error functions.
 We compare our algorithm with 3D-GS [17] and three state-of-the-art NeRF-based approaches: Mip-NeRF360 [2], InstantNGP [28] and Plenoxels [9].
-All methods were configured using the same settings as outlined in the 3D-GS paper to control variables, preventing the introduction of variables other than our Optimal Splatting method.
-6.4 Results
+All methods are configured using the same settings as outlined in the 3D-GS [17] to control variables, preventing the introduction of variables other than our Optimal Splatting method.
 Quantitative comparisons
@@
 Specifically, every 8th photo is reserved for testing, ensuring a consistent and meaningful basis for comparisons to generate error metrics.
-Standard metrics such as PSNR, LPIPS [43], and SSIM, widely used in the literature, are employed for evaluation; refer to Table 1 for details.
-It can be observed that, while the 3D-GS method significantly outperforms MipNeRF360 in rendering speed, its rendering quality slightly falls behind MipNeRF360, with a lower average PSNR across the 13 scenes.
+Standard metrics such as PSNR, LPIPS [45], and SSIM are employed for evaluation.
+We report quantiative results in Table 1.
+While the 3D-GS significantly outperforms MipNeRF360 in rendering performance, its rendering quality slightly falls behind MipNeRF360, with a lower average PSNR across the 13 scenes.
 This might be attributed to MipNeRF360 not employing significant approximations throughout its entire pipeline.
-However, by replacing the original projection method of 3D-GS with our Optimal Projection, which incurs smaller errors, we not only maintain rendering speed far surpassing MipNeRF360 but also achieve a quality improvement.
-Furthermore, in terms of average results across the 13 scenes, our approach outperforms others, including 3D-GS and several NeRF-based methods, in PSNR, SSIM, and LPIPS.
+However, by replacing the original projection method of 3D-GS with our optimal projection, which incurs smaller errors, we not only maintain rendering performance far surpassing MipNeRF360 but also achieve a quality improvement.
 Qualitative comparisons
-Due to space constraints, for our qualitative experiments, we selected only two methods, 3D-GS and MipNeRF360, from Table 1, which have comparable metrics, as our baselines.
-Please see Figure 5.
-From the comparative images, it is evident that our algorithm, compared to 3D-GS, reduces artifacts and defects while preserving more details.
-In the Truck scene, we address the absence of the steering wheel in the reconstruction by 3D-GS (indicated by the red arrow), while avoiding the pillar fragmentation introduced by MipNeRF360 (indicated by the blue arrow).
-Additionally, in the Train scene, we maintain the geometry of the train’s iron railing while avoiding the black floaters present in 3D-GS.
-In the PlayRoom scene, we achieve relatively better wire geometry.
-In the DrJohnson scene, our algorithm neither exhibits the fragmented appearance seen in 3D-GS nor the blurred details observed in MipNeRF360.
-In the Room and Bonsai scenes, we respectively have better stems in the flowerpot and improved details on the small white objects on the floor compared to 3D-GS.
-In the Counter scene, our floor texture is optimal.
-In the Stump scene, our tree branches do not fracture as observed in 3D-GS.
-7 Discussion and Conclusions
-Because Gaussian functions remain Gaussian after affine transformations, convolutions, and similar operations, but this is not necessarily the case for projection transformations, the 3D Gaussian Splatting process employs a Taylor first-order expansion to approximate this projection transformation as an affine transformation.
-However, such approximation inevitably introduces errors, which can result in artifacts, distortions, and inaccuracies in the learning comprehension of the scene as well as the generation of novel view images.
-This paper begins by examining the Taylor expansion remainder of the local affine approximation, obtaining the mathematical expectation of the error.
-This expectation is a function of the mean of the Gaussian function.
-By finding critical points, it is determined that the function attains its minimum when the projection of the mean of the Gaussian function on the plane coincides with the projection of the camera center on the plane.
-The visualization of this function is depicted in Figure 2.
-Based on the conclusions drawn from this theoretical analysis, we propose a more optimal projection method with reduced error, referred to as Optimal Gaussian Splatting.
-Subsequently, extensive experiments are conducted to further validate the superiority of this projection approach.
-The original 3D-GS [16], despite exhibiting significantly superior performance compared to the NeRF-based MipNeRF360 [2] method, may suffer a slight compromise in quality due to errors introduced by projection, as opposed to MipNeRF360 [2].
-In contrast, our Optimal Gaussian Splatting, by minimizing projection errors, unlocks a portion of the untapped potential of 3D-GS.
-This demonstrates that the representation of scenes using 3D-GS can achieve state-of-the-art (SOTA) levels in both quality and performance.
-Additionally, it is noteworthy that our entire error analysis process only considers the influence of Gaussian mean on the error function while treating the covariance as a constant.
-Hence, future research could explore optimizing 3D-GS further by considering the Gaussian covariance in the analysis.
-Furthermore, our method, throughout its entire process, is only dependent on the transformation from image pixels to the cutting plane (corresponding to Equation 23) and its relationship with the image representation.
-Unlike the original 3D-GS, where the $z=1$ plane tends to favor perspective images due to the perspective projection viewport transformation involving the NDC’s $z=1$ plane, our approach offers more versatility.
-This makes our method potentially applicable to various image representations.
-However, this versatility may come with a slight performance trade-off or require appropriate code implementation to avoid extending training time due to the additional operation of pixel-to-tangent-plane transformation, as shown in Equation 23.
-In summary, our work represents the first theoretical analysis that focuses on the projection error of 3D Gaussian Splatting.
-We have introduced a novel Splatting method Optimal Gaussian Splatting aimed at reducing projection errors, and through both theoretical analysis and empirical validation, demonstrated the superiority of this approach, as shown in Figure 1 and Figure 5.
+For our qualitative experiments, we select only two methods, 3D-GS and MipNeRF360, as our baselines.
+It can be observed that our method is capable of generating more realistic details, with fewer artifacts compared to 3D-GS [17].
+And this is precisely the superiority brought about by our projection method, which results in smaller errors.
+Furthermore, We see that in some cases Mip-NeRF360 has remaining artifacts that our method avoids (e.g., black lines - in Playroom).
+6.4 Discussions
+Impacts of Decreasing Focal Length
+Based on Equation 9, we further derive the projection error function with respect to image coordinates $(u,v)$ and focal length $f_{x},f_{y}$ through the coordinate transformation between image coordinates and polar coordinates $(\theta_{\mu},\phi_{\mu})$.
+As illustrated in Figure 5, the peak of projection error mainly occurs at the image edges under a long focal length, resulting in artifacts at the image edges.
+As the focal length decreases, the field of view expands, leading to more Gaussians deviating from the projection center and an overall increase in projection error.
+In such cases, 3D-GS exhibit more artifacts, such as needle-like structures or cloud-like large Gaussians, which obscure parts that perform well under long focal length, thereby significantly degrading the overall image quality.
+Since our method employs a central radial projection for the projection plane, theoretically, we should not encounter such defects.
+To validate this perspective, we conduct experiments at various focal length reduction ratios (reduction ratios of ×$\times 0.2$ and ×$\times 0.3$).
+Table 2: We quantitatively compare our method with the original 3D-GS [17] using focal lengths of 0.2 and 0.3 times the original.
+Six scenes were selected from three datasets: Mip-NeRF360 [2], Tank &\&& Templates [18], and Deep Blending [14].
+The scenes include Train and Truck from Tanks &\&& Templates, DrJohnson and Playroom from Deep Blending, and Room and Treehill from Mip-NeRF360.
+Metric Focal Method Train Truck Dr Johnson Playroom Room Treehill Average
+[…]
+Ours 16.60 16.70 23.54 23.29 23.67 18.98 20.46
+$\times$0.3 3D-GS [17] 12.85 14.89 23.36 22.51 25.00 18.35 19.49
+Ours 17.52 18.07 24.89 26.40 25.87 19.87 22.10
+[…]
+Ours 0.565 0.537 0.701 0.752 0.777 0.435 0.628
+$\times$0.3 3D-GS [17] 0.471 0.529 0.740 0.791 0.812 0.437 0.630
+Ours 0.588 0.605 0.762 0.844 0.834 0.472 0.684
+[…]
+Ours 0.277 0.250 0.216 0.203 0.178 0.385 0.251
+$\times$0.3 3D-GS [17] 0.331 0.276 0.207 0.213 0.154 0.397 0.263
+Ours 0.243 0.217 0.201 0.182 0.158 0.385 0.231
+Figure 6: We compare our method to the original 3D-GS [17] with different focal length reductions.
+Two focal lengths, 0.2 and 0.3 times the original, are tested.
+Ground truth ×$\times 1$ in the left images indicates no focal length reduction.
+The red-boxed areas in the right images correspond to the same regions in the left images.
+Figure 7: Renderings of three scenes using our projection under three camera models.
+From left to right: the ground truth of the original perspective, wide-angle rendering with a pinhole camera, fisheye camera rendering, and the panorama.
+For quantitative experiments, considering that there is no ground truth for large field-of-view in the 13 scenes from the 3D-GS [17], we apply a focal length mask.
+This mask selects central parts of the rendered images with corresponding ground truth based on the focal length scaling factor for metric calculation.
+The results are presented in the Table 2.
+Whether it is at ×$\times 0.2$ or ×$\times 0.3$ reduction ratios, our method outperforms the original 3D-GS significantly across various scenes in terms of all metrics.
+As the focal length decreases from ×$\times 0.3$ to ×$\times 0.2$, 3D-GS [17] experiences a significant degradation due to increased projection errors, while our method demonstrates greater robustness.
+For quantitative experiments, we randomly select three scenes, each with two different focal lengths for demonstration.
+Additional experimental results can be found in the supplementary materials.
+It is evident from the Figure 6 that our method exhibits greater robustness across various focal length settings, mitigating the artifacts introduced by projection in the original 3D-GS.
+Adaptability to Various Camera Models
+By straightforwardly modifying the transformation from image space to camera space, we have achieved adaptation to various camera models, with the caveat that the original 3D-GS’s projection struggles to support these camera models.
+Figure 7 illustrates that our projection, in addition to conforming to the original pinhole camera according to Equation 16, can also accommodate fisheye cameras and generate panoramas.
+Limitation
+Due to our method’s independence from the image plane, it exhibits adaptability to various camera models.
+However, the slight increase in training time is a consequence of the subsequent transformation to the image plane after projection onto the tangent plane, rather than a direct projection onto the image plane.
+Fortunately, this method avoids the need for dense points sampling, ensuring the algorithm’s time complexity remains unchanged and rendering performance stays unaffected.
+Moreover, it may be possible to mitigate the impact on training time, particularly concerning the corresponding backward propagation, through the implementation of optimized CUDA code.
+7 Conclusion
+Through theoretical analysis, we establish the correlation between the projection error and the Gaussian mean in 3D-GS.
+By optimizing the error function, we discover that the projection error is minimized when projecting on the plane perpendicular to the line connecting the Gaussian mean and the camera center.
+Leveraging this insight, we propose a novel projection that yields high-quality images without compromising rendering performance and can be easily adapted to various camera models through simple modifications.
+In this work, our primary focus is on analyzing the relationship between the projection error and the mean of the 3D Gaussian under the assumption of a constant covariance.
+The influence of Gaussian covariance on projection requires further discussion in future work.
+It may be possible to further explore the potential of Gaussian Splatting as an explicit representation technique for scenes.
+Appendix 0.A Details of Error Analysis
+Unit Vector Simplification
+We have proved that the composition of the transformation $\varpi$, projecting onto the unit sphere, and the transformation $\varphi$, projecting onto the tangent plane, is equivalent to the single transformation $\varphi$ directly projecting onto the tangent plane:
+$\begin{split}\left(\varphi\circ\varpi\right)\left({\mathbf{x}^{{}^{\prime}}}\right)&=\varphi\left({\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\right)\\ &={\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\left(\mathbf{x_{0}}^{\top}\left({\mathbf{x}^{{}^{\prime}}}{\left({\mathbf{x}^{{}^{\prime}}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1/2}}\right)\right)^{-1}\\ &={\mathbf{x}^{{}^{\prime}}}\left(\mathbf{x_{0}}^{\top}{\mathbf{x}^{{}^{\prime}}}\right)^{-1}\\ &=\varphi\left({\mathbf{x}^{{}^{\prime}}}\right)\text{.}\end{split}$ (19)
+Therefore, we can simplify the derivation of the error function $\epsilon$ by projecting all relevant points onto the unit sphere as unit vectors (Equation 8 in the main paper).
+Closed-Form Expression for the Error Function
+This section provides a comprehensive derivation of the integral expression for the error function as presented in the main paper.
+Upon expressing $\mathbf{x}^{{}^{\prime}}$, $\mathbf{x_{0}}$ and $\boldsymbol{\mu}^{{}^{\prime}}$ as unit vectors and substituting them into Equation 6 in the main paper, the Jacobian matrix is derived:
+$\mathbf{J}=\frac{\partial\varphi}{\partial\mathbf{x}^{{}^{\prime}}}\left(\boldsymbol{\mu}^{{}^{\prime}}\right)=\left[\begin{matrix}\frac{1}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}&0&-\frac{\sin{\left(\phi_{\mu}\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}\\ 0&\frac{1}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}&\frac{\sin{\left(\theta_{\mu}\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\theta_{\mu}\right)}}\\ 0&0&0\end{matrix}\right]\text{.}$ (20)
+Substituting this Jacobian matrix into Equation 5 in the main paper yields the derivation of the Taylor expansion remainder term:
+$R_{1}\left(\mathbf{x}^{{}^{\prime}}\right)=\left[\begin{matrix}-\frac{\sin{\left(\phi-\phi_{\mu}\right)}\cos{\left(\theta\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}+\tan{\left(\phi\right)}-\tan{\left(\phi_{\mu}\right)}\\ \frac{\sin{\left(\theta\right)}}{\cos{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}}-\frac{\sin{\left(\theta_{\mu}\right)}\cos{\left(\phi\right)}\cos{\left(\theta\right)}}{\cos^{2}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\theta_{\mu}\right)}}+\frac{\tan{\left(\theta_{\mu}\right)}}{\cos{\left(\phi_{\mu}\right)}}-\frac{\tan{\left(\theta\right)}}{\cos{\left(\phi\right)}}\\ 0\end{matrix}\right]\text{.}$ (21)
+And substituting this expression into Equation 7 in the main paper results in the integral expression (Equation 9 in the main paper).
+The main paper provides the integral expression (Equation 9 in the main paper) and the function graph (Figure 2 in the main paper) for the error function, without presenting the closed-form expression after integration.
+We provide the closed-form expression as follows:
+$\begin{split}\epsilon&\left(\theta_{\mu},\phi_{\mu}\right)=\frac{1}{{16\cos^{4}{\left(\phi_{\mu}\right)}\cos^{4}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}}}\\ &\Big{\{}8\left(\left(2\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}-\pi\right)\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+2\right)\cos^{4}{\left(\phi_{\mu}\right)}\cos^{4}{\left(\theta_{\mu}\right)}\tan^{2}{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+\\ &\left.\left(-4\cdot\left(2\cos{\left(2\theta_{\mu}\right)}+\pi\right)\sin^{2}{\left(\phi_{\mu}\right)}+\pi\left(2\cos{\left(2\theta_{\mu}\right)}+\pi\right)+4\cos{\left(2\theta_{\mu}\right)}+2\pi\right)\right.\\ &\left.\sin^{2}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+8\log{\frac{{\left(\sin{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+1\right)}{\left(\cos{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+1\right)}}{{\left(1-\sin{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}{\left(1-\cos{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}}}\right.\\ &\left.\log{\frac{{\left(1-\sin{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}{\left(1-\cos{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}{\exp{\left(2\sqrt{2}\cos{\left(\theta_{\mu}\right)}\right)}}}{{\left(\sin{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+1\right)}{\left(\cos{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+1\right)}{\left(\tan^{2\sin{\left(\theta_{\mu}\right)}}{\left(\theta_{\mu}+\frac{\pi}{4}\right)}\right)}}}\right.\\ &\left.\cos^{3}{\left(\phi_{\mu}\right)}\cos^{3}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+\right.\\ &\left.4\pi\left(-4\log{\left(\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}\tan{\left(\phi_{\mu}\right)}+\pi\tan^{2}{\left(\phi_{\mu}\right)}+2\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}-\pi\right)\right.\\ &\left.\cos^{4}{\left(\phi_{\mu}\right)}\cos^{4}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+\right.\\ &\left.\left(2\pi\left(\left(4+2\pi+16\sqrt{2}\right)\sin^{2}{\left(\theta_{\mu}\right)}-2+\pi\right)\cos^{2}{\left(\phi_{\mu}\right)}+\left(2\pi-4\right)\cos{\left(2\theta_{\mu}\right)}-2\pi+\pi^{2}\right)\right.\\ &\left.\cos^{2}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+\right.\\ &\left.16\sqrt{2}\log{\frac{{\left(1-\sin{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}{\left(1-\cos{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\right)}{\exp{\left(2\sqrt{2}\cos{\left(\theta_{\mu}\right)}\right)}}}{{\left(\sin{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+1\right)}{\left(\cos{\left(\phi_{\mu}+\frac{\pi}{4}\right)}+1\right)}}}\right.\\ &\left.\cos^{3}{\left(\phi_{\mu}\right)}\cos^{4}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+\right.\\ &\left.64\sin^{2}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\phi_{\mu}\right)}\cos^{4}{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}-\right.\\ &\left.16\sqrt{2}\cdot\left(1+\sqrt{2}\right)\sin{\left(\theta_{\mu}\right)}\sin{\left(2\theta_{\mu}\right)}\cos^{2}{\left(\phi_{\mu}\right)}\cos{\left(\theta_{\mu}\right)}\tan{\left(\phi_{\mu}+\frac{\pi}{4}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+\right.\\ &8\pi\cos^{4}{\left(\phi_{\mu}\right)}\cos^{4}{\left(\theta_{\mu}\right)}\tan{\left(\theta_{\mu}+\frac{\pi}{4}\right)}+\frac{8\left(-2\pi\cos^{2}{\left(\theta_{\mu}\right)}+\pi+4\right)\cos^{4}{\left(\phi_{\mu}\right)}\cos^{2}{\left(\theta_{\mu}\right)}}{\left(\tan{\left(\theta_{\mu}\right)}-1\right)^{2}}\Big{\}}\text{.}\end{split}$ (22)
+Appendix 0.B Details of Adaptation for Various Camera Models
+In the main paper, it is highlighted that our projection, independent of perspective image plane, allows for the adaptation to various camera models by modifying the transformation from image space to camera space in the rasterization based on the unit sphere (Equation 16 in the main paper) .
+This section provides a detailed explanation of our method’s adaptation to various camera models.
+Fisheye
+The design models for fisheye cameras can generally be categorized into four types: equidistant projection model, equisolid angle projection model, orthographic projection model, and stereographic projection model.
+We explane our projection’s adaptation to fisheye camera models, taking the equidistant projection model as an example.
+For a pixel $\left(u,v\right)$ on the image, we cast a ray. According to the transformation between the image space and camera space for the equidistant projection model, we obtain:
+$\mathbf{x}_{\text{2D}}=\varphi_{\text{p}}\left(\left[\begin{matrix}\frac{\left(u-c_{x}\right)\sin{\left(\sqrt{\frac{\left(u-c_{x}\right)^{2}}{f_{x}^{2}}+\frac{\left(v-c_{y}\right)^{2}}{f_{y}^{2}}}\right)}}{f_{x}\sqrt{\frac{\left(u-c_{x}\right)^{2}}{f_{x}^{2}}+\frac{\left(v-c_{y}\right)^{2}}{f_{y}^{2}}}}\\ \frac{\left(v-c_{y}\right)\sin{\left(\sqrt{\frac{\left(u-c_{x}\right)^{2}}{f_{x}^{2}}+\frac{\left(v-c_{y}\right)^{2}}{f_{y}^{2}}}\right)}}{f_{y}\sqrt{\frac{\left(u-c_{x}\right)^{2}}{f_{x}^{2}}+\frac{\left(v-c_{y}\right)^{2}}{f_{y}^{2}}}}\\ \cos{\left(\sqrt{\frac{\left(u-c_{x}\right)^{2}}{f_{x}^{2}}+\frac{\left(v-c_{y}\right)^{2}}{f_{y}^{2}}}\right)}\end{matrix}\right]\right)$ (23)
+where $c_{x}$, $c_{y}$, $f_{x}$, $f_{y}$ denote the intrinsic parameters of the camera model.
+Panorama
+For a pixel $\left(u,v\right)$ on the image, we cast a ray. According to the transformation between the image space and camera space for panoramic images, we obtain:
+$\mathbf{x}_{\text{2D}}=\varphi_{\text{p}}\left(\left[\begin{matrix}\sin{\left(\frac{\pi\left(-W+2u\right)}{W}\right)}\cos{\left(\frac{\pi\left(-\frac{H}{2}+v\right)}{H}\right)}\\ \sin{\left(\frac{\pi\left(-\frac{H}{2}+v\right)}{H}\right)}\\ \cos{\left(\frac{\pi\left(-\frac{H}{2}+v\right)}{H}\right)}\cos{\left(\frac{\pi\left(-W+2u\right)}{W}\right)}\end{matrix}\right]\right)$ (24)
+where $H,W$ represent the height and width of the image, respectively.
+Then, similar to Equation 17 in the main paper, the 2D Gaussian function values can be obtained for alpha blending to generate the image.
+The alpha blending process after obtaining the 2D Gaussian is the same as in the original 3D-GS [17].
+Figure 8: We show comparisons of our method to the original 3D-GS [17] under various camera models and different focal lengths.
+Appendix 0.C Details of Experiments
+Additional Results
+Table 3-4 list PSNR for our evaluation over all considered techniques and real-world scenes, corresponding to Table 1 in the main paper.
+Figure 9 illustrates additional qualitative comparisons between our method and other approaches.
+Figure 9: We show comparisons of our method to previous methods and the corresponding ground truth images from held-out test views.
+The scenes are, from the top down: Train from Tanks&Temples [18]; Drjohnson from the Deep Blending dataset [14] and Room, Stump from Mip-NeRF360 dataset [2].
+Differences in quality highlighted by arrows/insets.
+Impacts of Decreasing Focal Length
+Figure 10: Illustration of details of metrics caculation in the quantitative comparisons on the impact of decreasing focal length.
+In the quantitative comparison experiments about the impact of reducing focal length, we utilized a focal length mask methodology due to the absence of ground truth for wide-angle images.
+The detailed procedures are illustrated in Figure 10.
+The focal mask selects central parts of the rendering with corresponding ground truth based on the focal length scaling factor for metric calculation.
+For rendering, we crop the corresponding part with available ground truth, then resize the ground truth to match the size of this patch.
+Finally, we compute the metrics for these two images.
+For qualitative experiments, we show additional results in Figure 8.
+In addition to the original focal length reduction factors of 0.3 and 0.2, we introduce 0.1 times the focal length.
+Additionally, we generate images with corresponding perspectives using other camera models, while 3D-GS [17] is not presented as it lacks support for these camera models.
+Comparison with the Other 3D-GS Methods
+We choose Mip-Splatting [44] and Scaffold-GS [22] to compare in the same setting as in Figure 6 of the main paper.
+For Mip-Splatting [44] and Scaffold-GS [22] still employ conventional projection of 3D-GS, the approximation error remains larger compared to ours, as shown in Figure 11.
+Additionally, our method can be be seamlessly integrated with theirs to reduce the projection error of these methods.
+Figure 11: We show comparisons of our method to the original 3D-GS [17], Mip-Splatting [44] and Scaffold-GS [22] with a large FOV camera.
+In the figure, Ours(+X) indicates the combination of our method with the X method.
+Adaptability to Various Camera Models
+Since our method is based on differentiable rasterizer, it can also be used to train on the non-pinhole camera dataset.
+We illustrate the results obtained from training on the non-pinhole camera dataset Matterport [5] directly, where 3D-GS fails entirely to reconstruct the scene, as shown in Figure 12.
+Figure 12: We show the results of our method trained on the non-pinhole camera dataset Matterport [5].
+Table 3: PSNR scores for Mip-NeRF360 scenes.
+MipNeRF360 Indoor MipNeRF360 Outdoor Avg.
+Room Counter Kitchen Bonsai Avg. Bicycle Flowers Garden Stump Treehill Avg.
+Plenoxels [9] 27.59 23.62 23.42 24.67 24.83 21.91 20.10 23.49 20.66 22.25 21.68 23.08
+INGP-Base [28] 29.27 26.44 28.55 30.34 28.65 22.19 20.35 24.60 23.63 22.36 22.63 25.30
+INGP-Big [28] 29.69 26.69 29.48 30.69 29.14 22.17 20.65 25.07 23.47 22.37 22.75 25.59
+M-NeRF360 [2] 31.63 29.55 32.23 33.46 31.72 24.37 21.73 26.98 26.40 22.87 24.47 27.69
+3D-GS [17] 30.63 28.70 30.32 31.98 30.41 25.25 21.52 27.41 26.55 22.49 24.64 27.21
+Ours 31.58 29.05 31.23 32.31 31.04 25.07 21.54 27.15 26.57 22.77 24.62 27.48
+Table 4: PSNR scores for Tanks&Temples and Deep Blending scenes.
+Tanks&Temples Deep Blending Avg.
+Truck Train Avg. Dr Johnson Playroom Avg.
+Plenoxels [9] 23.22 18.93 21.07 23.14 22.98 23.06 22.07
+INGP-Base [28] 23.26 20.17 21.72 27.75 19.48 23.62 22.67
+INGP-Big [28] 23.38 20.46 21.92 28.26 21.67 24.96 23.44
+M-NeRF360 [2] 24.91 19.52 22.22 29.14 29.66 29.40 25.81
+3D-GS [17] 25.19 21.10 23.14 28.77 30.04 29.41 26.27
+Ours 25.19 21.67 23.43 29.14 29.89 29.51 26.47